### v3

# 7-3-1 Hongo, Bunkyo-ku, Tokyo 113-0033, Japan

Abstract
We apply the background charged bosonic free-field approach to the rational principal quantum Drinfeld-Sokolov (QDS) $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ minimal models with boundaries, where $g$ is a simple bosonic Lie algebra. We express their Ishibashi states using the free-field Ishibashi states, and calculate disk two-point correlation functions using the free-field vertex operators with intertwiner insertions. The integral results are obtained by repeatedly applying the Euler-type integral expression and the Taylor expansions of Lauricella’s hypergeometric functions $F_{D}^{(n)}$. We also discuss the background free-field approach to the QDS $\mathcal{W}$ algebras related to a semi-simple Lie algebra $g$, and apply a coset free-field resolution to express diagonal ADE coset $\mathcal{W}$ minimal Ishibashi states.
1 Introduction
Two-dimensional conformal field theories (CFT2) have infinite-dimensional local conformal symmetries, described by a local conformal algebra $\mathcal{A}\otimes\widebar{\mathcal{A}}$. $\mathcal{A}$ and $\widebar{\mathcal{A}}$ are infinite-dimensional Lie algebras, called the chiral and anti-chiral algebras, respectively Belavin:1984vu ; Moore:1988qv . The infinite-dimensional local conformal symmetries allow us to solve the correlation functions in some theories exactly, and free-field approaches are efficient methods. We claim that a CFT2 admits a free-field approach if the following three conditions are satisfied Bouwknegt:1990wa .
  • $\mathcal{A}$ and $\widebar{\mathcal{A}}$ can be realized using the free-fields.
  • The existence of projection maps from the free-field Fock space modules to chiral modules, which are referred to as the free-field resolutions or Fock space resolutions.
  • The existence of a method of calculating the correlation functions using the free-field vertex operators and the free-field $\mathcal{A}$ intertwiners.
There is a wide range of RCFTs that are known to admit free-field approaches. Two canonical types of theories with free-field approaches are
We apply the background-charged free-field approach to rational QDS $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ ($g$ being simple Lie algebras), including free-field expressions of Ishibashi states, and detailed example calculations of disk correlation functions using the free-field vertex operators, with insertions of $\mathcal{W}$ intertwiners. The method to calculate the correlation functions is a direct generalization of the celebrated Coulomb gas formalism of the Virasoro minimal models Dotsenko:1984nm ; Dotsenko:1984ad ; Dotsenko:1985hi . We show that the integral expressions of this type are solutions to the Belavin-Polyakov-Zamolodchikov (BPZ) type of differential equations of the QDS $\mathcal{W}$ minimal models Belavin:1984vu . The crucial techniques in performing the integrals obtained from the free-field approaches are:
  • Deforming the contour integrals into a chained integral along real segments.
  • The application of the Euler-type integral expression and the Taylor expansions of Lauricella’s hypergeometric functions $F_{D}^{(n)}$.
Divergent terms can arise in the process. However, we have shown that the integrals expressions are solutions to the BPZ type of differential equations, it is necessary to include them. We give sample calculations of disk two-point functions in some diagonal and charge-conjugated Virasoro minimal models, $\mathcal{W}_{3}$ minimal models, the charge-conjugated $\mathcal{W}_{4}(7,4)$ model, and the $\mathcal{W}\big{[}\widehat{G}_{2}\big{]}(6,7)$ model. The calculations of the correlation functions with higher points and in other rational $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ models can be obtained from analogous procedures in principle.
In the appendices, we discuss the background-charged free-field approach to the QDS $\mathcal{W}$ algebras related to semi-simple Lie algebras. The free-field approach reveals that each simple part of a semi-simple QDS $\mathcal{W}$ algebra decouples. The chiral characters, modular properties, fusion rules, and correlation functions of the semi-simple $\mathcal{W}$ theory are a multiplication of those of the simple parts. Finally, we discuss applying a coset resolution Bouwknegt:1990fb ; Bouwknegt:1990wa to express the diagonal ADE coset $\mathcal{W}$ minimal Ishibashi states Goddard:1984vk ; Goddard:1986ee ; Mathieu:1990dy ; Bouwknegt:1992wg ; Arakawa:2018iyk .
2 Background-charged bosonic field approach to the principal quantum Drinfeld-Sokolov $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ minimal model with boundaries
2.1 Preliminaries
Concepts in simple Lie and corresponding KM algebras.
A finite Lie algebra $g$ has $r$ Abelian generators $H^{i}$, where $r$ is called its rank. The rest $g$ generators are called ladder operators $E^{\alpha}$, where each of them is related to an $r$-dimensional vector $\alpha$ called the roots, whose set is denoted by $\Delta$. The roots with the first nonzero element being positive are called positive roots, whose set is denoted by $\Delta_{+}$. For each root $\alpha$, we define its corresponding coroot as $\alpha^{\vee}=\frac{2\alpha}{|\alpha|^{2}}$. There are a total of $r$ simple roots, denoted by $\alpha_{i}$. The Cartan matrix of $g$ is defined as
$A_{ij}=(\alpha_{i},\alpha_{j}^{\vee}).$ (1)
We consider simple Lie algebras $g$, which include four infinite series $A_{r\geq 1}\cong sl(r+1;\mathbb{C})$, $B_{r\geq 2}=so(2r+1;\mathbb{C})$, $C_{r\geq 2}\cong sp(2r;\mathbb{C})$, $D_{r\geq 3}\cong so(2r;\mathbb{C})$, and five exceptional algebras $E_{6,7,8}$, $F_{4}$, and $G_{2}$. For a simple Lie algebra $g$, there exists a unique highest root $\theta$, where the summation of linear coefficients of simple root expansions is maximized. The linear coefficients of the $\theta$ expansion, in terms of $\alpha_{i}$ and $\alpha_{i}^{\vee}$, are called the marks $a_{i}$ and the comarks $a_{i}^{\vee}$ respectively
$\theta=\sum_{i=1}^{r}a_{i}\alpha_{i}=\sum_{i=1}^{r}a_{i}^{\vee}\alpha_{i}^{\vee}.$ (2)
We define a parameter $r^{\vee}$ as
$r^{\vee}:=\text{max}\big{\{}a_{i}/a_{i}^{\vee}\big{\}}.$ (3)
If $r^{\vee}=1$, the simple Lie algebra is called simply-laced. The $A_{r\geq 1}$, $D_{r\geq 3}$, and $E_{6,7,8}$ algebras are simply-laced. The $B_{r\geq 2}$, $C_{r\geq 2}$ series, and the $F_{4}$ algebra have $r^{\vee}=2$. The $G_{2}$ algebra has $r^{\vee}=3$. A finite Lie algebra $g$ is called semi-simple if it is a finite direct sum of simple Lie algebras $g=g_{1}\oplus g_{2}\oplus\cdots\oplus g_{n}$.
Weight vectors of $g$ are eigen-vectors of the Abelian generators. We express the $g$ weight vectors $\lambda$ using the Dynkin labels $\lambda_{i}$
$\lambda=\sum_{i=1}^{r}\lambda_{i}\omega_{i}=[\lambda_{1},\lambda_{2},\cdots,\lambda_{r}],\quad(\omega_{i},\alpha_{j}^{\vee})=\delta_{ij}$ (4)
where $\omega_{i}$ are the fundamental weights of $g$. We define the Weyl vector $\rho$ and the dual Weyl vector $\rho^{\vee}$ of $g$ by $(\rho,\alpha_{i}^{\vee})=1$ and $(\rho^{\vee},\alpha_{i})=1$, $\forall i$. Express them in terms of Dynkin labels
$\rho=[1,\cdots,1],\quad\rho^{\vee}=\Big{[}\frac{a_{1}}{a_{1}^{\vee}},\cdots,\frac{a_{i}}{a_{i}^{\vee}}\Big{]}.$ (5)
A weight vector $\lambda$ is called dominant if $\lambda_{i}\in\mathbb{N}$, with at least one positive integer Dynkin label. The set of dominant $g$ weight vectors is denoted by $P_{+}$. A codominant weight vector can be defined using relation $(\lambda,\alpha_{i})\in\mathbb{N}$, with at least one positive integer inner product. The set of codominant weight vectors is denoted by $P_{+}^{\vee}$. The weight lattice $P$, root lattice $Q$, and coroot lattice $Q^{\vee}$ of $g$ are defined as
$P=\sum_{i=1}^{r}\mathbb{Z}\omega_{i},\quad Q=\sum_{i=1}^{r}\mathbb{Z}\alpha_{i},\quad Q^{\vee}=\sum_{i=1}^{r}\mathbb{Z}\alpha_{i}^{\vee}.$ (6)
The number of elements in their cosets can be calculated as $|P/Q^{\vee}|=\det(\alpha_{i}^{\vee},\alpha_{j}^{\vee})$ and $|P/Q|=\det(A_{ij})$.
The Weyl group $W(g)$ of a bosonic Lie algebra $g$ is a finite-dimensional group, generated by the simple Weyl reflections $w_{i}$, $i=1,\cdots,r$
$w_{i}\lambda:=\lambda-\lambda_{i}\alpha_{i}.$ (7)
For each element $w\in W(g)$, we can define its length $l(w)$ as the minimum possible number of elements in its simple Weyl reflection decompositions. There is a unique longest element $w_{l}\in W(g)$ and the charge conjugation of $\lambda$ is defined as $U\lambda:=-w_{l}\lambda$. For the $B_{r}$, $C_{r}$, $D_{r}$ even series, $E_{7,8}$, $F_{4}$, and $G_{2}$ the charge conjugation $U=-w_{l}=I$. For the $A_{r}$ series, $U$ reverse the order of all Dynkin labels. For the $D_{r}$ odd series, $U$ reverse the final two Dynkin labels $\lambda_{r-1}\leftrightarrow\lambda_{r}$. For $E_{6}$, $U$ reverse two pairs of Dynkin labels $\lambda_{1}\leftrightarrow\lambda_{6}$ and $\lambda_{3}\leftrightarrow\lambda_{5}$. We define two types of shifted Weyl actions as
$w\ast\lambda:=w(\lambda+\rho)-\rho,\quad w\ast^{\vee}\lambda:=w(\lambda+\rho^{\vee})-\rho^{\vee}.$ (8)
Now, we extend the concepts into the KM algebras related to a simple Lie algebra $\widehat{g}(k)$ (which will be called simple KM algebras) Kac:1984mq . A simple KM algebra is a $\mathbb{Z}_{n}$-graded extension of $g$, whose generators are
$H^{i}_{n}:=H^{i}\otimes z^{n},\quad E^{\alpha}_{n}:=E^{\alpha}\otimes z^{n},\quad z\in\mathbb{C}.$ (9)
Define the Coxeter number $h$ and the dual Coxeter numbers $h^{\vee}$ of $\widehat{g}(k)$ as
$h=1+\sum_{i=1}^{r}a_{i},\quad h^{\vee}=1+\sum_{i=1}^{r}a^{\vee}_{i}.$ (10)
A simple KM algebra has a unique central extension called the level $k$. The central extension extends the KM weight vectors $\widehat{\lambda}$ to $(r+1)$-dimensional
$\widehat{\lambda}=[\lambda_{0},\cdots,\lambda_{r}],\quad\widehat{\lambda}_{0}=k-\sum_{i=1}^{r}a^{\vee}_{i}\lambda_{i}.$
Dominant and codominant weight vectors are defined analogously to the finite cases, with the addition of the extra zeroth dimension. The set of level-$k$ dominant and codominant $\widehat{g}(k)$ weight vectors are denoted by $\widehat{P}_{+}^{k}$ and $\widehat{P}_{+}^{\vee k}$ respectively. The principal admissible weight vectors are defined in Frenkel:1992ju . We are only interested in the admissible $\widehat{g}(k)$ weights taking the form of
$\widehat{\Lambda}=\alpha_{+}\widehat{\Lambda}^{(+)}+\alpha_{-}\widehat{\Lambda}^{(-)},\quad\widehat{\Lambda}^{(+)}\in\widehat{P}_{+},\quad\widehat{\Lambda}^{(-)}\in\widehat{P}_{+}^{\vee}$. (11)
We call their finite part as admissible $g$-weight vectors, denoted by $\Lambda=(\Lambda^{(+)},\Lambda^{(-)})=\alpha_{+}\Lambda^{(+)}+\alpha_{-}\Lambda^{(-)}$. The parameters $\alpha_{\pm}$ are defined by $\alpha_{+}^{2}=(k+h^{\vee})^{-1}$ and $\alpha_{+}\alpha_{-}=-1$.
[…]
The free-field resolution conjectures of dominant and admissible $\widehat{g}(k)$ highest-weight modules.
Define the screening operators of a chiral algebra $\mathcal{A}$ as the free-field vertex operators whose OPEs with free-field chiral generators are either a regular or a total derivative singular term. From the definition, we conclude that a closed contour integral over an $\mathcal{A}$ screening operator is an $\mathcal{A}$ intertwiner.
The free-field approach to the simple KM $\widehat{g}(k)$ algebras we used here is the Wakimoto-type, which is realized from $r$ background-charged bosonic fields $\phi^{i}$ and $\big{|}\Delta_{+}\big{|}$ bosonic $\beta\gamma$ ghost systems Wakimoto:1986gf ; Bouwknegt:1990aa ; Bouwknegt:1990wa . Our focus is to present a free-field resolution conjecture for dominant and admissible modules $\widetilde{L}_{\widehat{\Lambda}}$ using the $\phi^{i}\beta^{\alpha}\gamma^{\alpha}$ Fock space modules, since the Fock space resolution of irreducible QDS $\mathcal{W}$ minimal modules are induced from them. Consider a simple KM algebra at positive integer level $k\in\mathbb{Z}_{+}$, where the chiral modules are the irreducible dominant modules $\widetilde{L}_{\widehat{\Lambda}}$, $\widehat{\Lambda}\in\widehat{P}^{k}_{+}$. The screening operators of this theory take the form of Bouwknegt:1990wa
$\widetilde{s}^{+}_{i}(z)=\>:e^{-i\alpha_{+}\alpha_{i}\cdot\phi}:(z)\>\widetilde{P}$, (12)
where $\widetilde{P}$ are polynomials of the $\beta\gamma$ fields.
[…]
$\quad\cdots\longrightarrow\widetilde{F}_{\widehat{\Lambda}}^{(-2)}\overset{\widetilde{d}^{-1}}{\longrightarrow}\widetilde{F}_{\widehat{\Lambda}}^{(-1)}\overset{\widetilde{d}^{0}}{\longrightarrow}\widetilde{F}_{\widehat{\Lambda}}^{(0)}\overset{\widetilde{d}^{1}}{\longrightarrow}\widetilde{F}_{\widehat{\Lambda}}^{(1)}\longrightarrow\cdots,$ (13)
where $\widetilde{F}_{\widehat{\Lambda}}=F_{\widehat{\Lambda}}^{\phi^{i}}\otimes F^{\beta^{\alpha}\gamma^{\alpha}}$, and $\widetilde{F}_{\widehat{\Lambda}}^{(n)}$ is
$\widetilde{F}_{\widehat{\Lambda}}^{(n)}=\oplus_{\widehat{w}}^{l(\widehat{w})=n}\Big{[}F^{\phi^{i}}_{\widehat{w}\ast\widehat{\Lambda}}\otimes F^{\beta^{\alpha}\gamma^{\alpha}}\Big{]}.$ (14)
$\widetilde{d}$ are the $\widehat{g}(k)$ intertwiners that shift the $\widehat{g}(k)$ weight by the required amount. It is conjectured that the zeroth cohomology space of the complex $\widetilde{C}_{\widehat{\Lambda}}$ is isomorphic to the dominant module $\widetilde{L}_{\widehat{\Lambda}}$ Bouwknegt:1989xa ; Bouwknegt:1989jf ; Bouwknegt:1990wa
$H^{0}\big{[}\widetilde{C}_{\widehat{\Lambda}}\big{]}\quad\cong\quad\widetilde{L}_{\widehat{\Lambda}}$. (15)
For the $\widehat{A}_{1}$ cases, the above conjecture is proven Bernard:1989iy . The dominant resolution conjecture is extended to the admissible modules $\widetilde{L}_{(\widehat{\Lambda}^{(+)},\widehat{\Lambda}^{(-)})}$ Bouwknegt:1991gf
$\widehat{\Lambda}=\alpha_{+}\widehat{\Lambda}^{(+)}+\alpha_{-}\widehat{\Lambda}^{(-)},\quad\widehat{\Lambda}^{(+)}\in\widehat{\Lambda}^{(+)}\in\widehat{P}_{+}^{p-h^{\vee}},\quad\widehat{\Lambda}^{(-)}\in\widehat{P}_{+}^{\vee\>p^{\prime}-h},$
$\alpha_{+}^{2}=\frac{1}{k+h^{\vee}}=\frac{p^{\prime}}{p},\quad\gcd(p,p^{\prime})=1,\quad\gcd(p^{\prime},r^{\vee})=1,\quad p\geq h^{\vee},\quad p^{\prime}\geq h.$ (16)
The resolution complexes $\widetilde{C}_{(\widehat{\Lambda}^{(+)},\widehat{\Lambda}^{(-)})}$ for the admissible modules are obtained by replacing the Fock space modules in the dominant resolution complexes by
$\widetilde{F}^{(n)}_{(\widehat{\Lambda}^{(+)},\widehat{\Lambda}^{(-)})}=\oplus_{\widehat{w}}^{l(\widehat{w})=n}\Big{[}F^{\phi^{i}}_{(\widehat{w}\ast\widehat{\Lambda}^{(+)},\widehat{\Lambda}^{(-)})}\otimes F^{\beta^{\alpha}\gamma^{\alpha}}\Big{]},$ (17)
and the admissible module $\widetilde{L}_{(\widehat{\Lambda}^{(+)},\widehat{\Lambda}^{(-)})}$ is conjectured to be isomorphic to the zeroth cohomology space of the complex $\widetilde{C}_{(\widehat{\Lambda}^{(+)},\widehat{\Lambda}^{(-)})}$.
Background-charged free boson approach to principal QDS $\mathcal{W}$ algebras.
The quantum Drinfeld-Sokolov QDS $\mathcal{W}$ algebras, related to simple Lie algebras $g$, are obtained from the following logic: we begin with a simple KM algebra $\widehat{g}(k)$, imposing a first-class constraint condition on it, and define the QDS $\mathcal{W}$ algebra as the BRST cohomology algebra Bouwknegt:1989xa ; Bouwknegt:1990wa ; Bouwknegt:1992wg ; Kac:2003jh . The $\mathcal{W}$ algebras considered in this work are obtained from imposing the principal embedding condition, and they are denoted by $\mathcal{W}\big{[}\widehat{g}(k)\big{]}$. The QDS $\mathcal{W}$ algebras obtained by imposing different constraint conditions exceeds the scale of this work.
The following lemma $H_{Q}\big{(}F_{\beta\gamma}\otimes F_{\text{bc}}\big{)}\cong\mathbb{C}$ (proven in Bershadsky:1989mf ) indicates that the $r$ background-charged fields $\phi^{i}$ provides a natural free-field approach for the theories with the QDS $\mathcal{W}$ algebra. The free-field $\mathcal{W}\big{[}\widehat{g}(k)\big{]}$ energy-stress tensor is given by Bouwknegt:1992wg
$T(z)=-\frac{1}{2}:\partial\phi\cdot\partial\phi:(z)-(\alpha_{+}\rho+\alpha_{-}\rho^{\vee})\cdot i\partial^{2}\phi(z).$ (18)
The $\mathcal{W}\big{[}\widehat{g}(k)\big{]}$ screening operators are
$s^{\pm}_{i}(z)=\>:e^{-i\alpha_{\pm}\alpha_{i}\cdot\phi}:(z).$ (19)
Let $C_{(\Lambda^{(+)},\Lambda^{(-)})}=(F_{(\Lambda^{(+)},\Lambda^{(-)})}^{\phi^{i}},d)$ be the BRST image of the $\widehat{g}(k)$ admissible resolution complex $\widetilde{C}_{(\widehat{\Lambda}^{(+)},\widehat{\Lambda}^{(-)})}$, where $d$ are $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ intertwiners. The BRST image of an admissible (and dominant) $\widehat{g}(k)$ modules is either zero or irreducible modules $L_{(\Lambda^{(+)},\Lambda^{(-)})}$ of the $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ algebras Frenkel:1992ju . Hence, the zeroth cohomology of the complex $(F_{(\Lambda^{(+)},\Lambda^{(-)})},d)$ is either trivial or being isomorphic to the irreducible modules $L_{(\Lambda^{(+)},\Lambda^{(-)})}$ Bouwknegt:1991gf .
Rational $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ minimal models.
We consider rational $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ minimal models, which are constrained by the torus partition function modular invariant and the finite closed fusion algebra conditions Cardy:1986ie ; Moore:1988qv . The Kac-tables of the rational $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ minimal models are Bouwknegt:1992wg
$\Big{\{}L_{(\Lambda^{(+)},\Lambda^{(-)})}\quad|\quad\widehat{\Lambda}^{(+)}\in\widehat{P}^{p-h^{\vee}}_{+},\>\widehat{\Lambda}^{(-)}\in\widehat{P}^{p^{\prime}-h}_{+}\Big{\}},$ (20)
with a Weyl symmetry identification
$\widehat{\Lambda}^{(+)}\to w\ast\widehat{\Lambda}^{(+)}+p\beta,\quad\widehat{\Lambda}^{(-)}\to w\ast^{\vee}\widehat{\Lambda}^{(-)}+p^{\prime}\beta,\quad\beta\in Q^{\vee}_{L},$ (21)
where $Q^{\vee}_{L}$ is the long root lattice of $g$. The $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ modular $S$ matrices take the form of Lukyanov:1990tf ; Mathieu:1990dy ; Frenkel:1992ju ; Arakawa:2016lef
$S_{(\Lambda^{(+)},\Lambda^{(-)})(\mu^{(+)},\mu^{(-)})}=\frac{1}{\sqrt{(pp^{\prime})^{r}}}\Big{|}\frac{P}{Q^{\vee}}\Big{|}^{-\frac{1}{2}}e^{2\pi i(\Lambda^{(+)}\mu^{(-)}+\Lambda^{(-)}\mu^{(+)})}$
$\sum_{w\in W(g)}\epsilon(w)e^{-\frac{2\pi ip^{\prime}}{p}\big{[}(\Lambda^{(+)}+\rho),\>w(\mu^{(+)}+\rho)\big{]}}\sum_{w^{\prime}\in W(g)}\epsilon(w^{\prime})e^{-\frac{2\pi ip}{p^{\prime}}\big{[}(\Lambda^{(-)}+\rho),\>w^{\prime}(\mu^{(-)}+\rho)\big{]}}.$ (22)
The fusion rules of RCFTs can be obtained from the Verlinde formula Dijkgraaf:1988tf ; Verlinde:1988sn
$N_{ij}^{k}=\sum_{l}\frac{S_{il}S_{jl}S_{kl}^{-1}}{S_{0l}}.$ (23)
The fusion rules related to the identity module are called obvious rules in this work.
Cardy boundary states.
We consider disk correlation functions with Cardy boundary conditions, hence we review the concept of Cardy boundary states Cardy:1989ir . Consider a CFT2 defined on the upper-half plane (UHP), the chiral generators satisfy the gluing conditions on the boundary (only the chiral symmetry preserved cases are considered) Cardy:1984bb ; Recknagel:1997sb ; Recknagel:1998ih
$T(z)=\widebar{T}(\widebar{z}),\quad W^{i}(z)=\Omega\widebar{W}^{i}(\widebar{z}),$ (24)
where $\Omega$ is an automorphism of the chiral algebra $\mathcal{A}$. Mapping the UHP to the unit disk, and assigning a boundary state $|b\rangle_{\Omega}$ to its boundary, we obtain the conditions
$(L_{n}-\widebar{L}_{-n})|b\rangle_{\Omega}=\big{[}W_{n}^{i}-\Omega(-1)^{h_{i}}\widebar{W}^{i}_{-n}\big{]}|b\rangle_{\Omega}=0.$ (25)
For each $\mathcal{A}$ representation $L_{i}$ (not necessarily irreducible), the solutions to the conditions (25) are the $\Omega$-twisted Ishibashi states $|i,b\rangle\rangle_{\Omega}$ Ishibashi:1988kg ; Onogi:1988qk
$|i\rangle\rangle_{\Omega}=\sum_{N=0}^{\infty}|i,N\rangle\otimes V_{\Omega}U|i,N\rangle,$ (26)
where $|i,N\rangle$ denotes the orthonormal basis of $L_{i}$. $U$ is an anti-unitary isomorphism mapping a representation to its charge conjugation. $V_{\Omega}$ is a unitary isomorphism induced from $\Omega$ such that $V_{\Omega}:\>L_{i}\to L_{\omega^{-1}(i)}$ Recknagel:1997sb ; Recknagel:2013uja . The physical boundary states need to be constrained by the following condition: the open-sector Hilbert space between two boundaries must take the form of
$\mathcal{H}^{b_{1}b_{2}}=\bigoplus_{i}\>n_{b_{1}b_{2}}^{i}L_{i},\quad n_{b_{1}b_{2}}^{i}\in\mathbb{N}.$ (27)
One simple type of solution satisfying this constraint is the Cardy boundary states, where the boundary states have the same labeling as the modules in the Kac-table, with the celebrated $n=N$ relation is imposed Cardy:1989ir . The forms of Cardy boundary states are
$|b_{i}\rangle_{\Omega}=\sum_{j}\frac{S_{ij}}{\sqrt{S_{0j}}}|j\rangle\rangle_{\Omega}.$ (28)
2.2 Universal results
Free-field expressions of $\mathcal{W}$ minimal Ishibashi states.
Consider diagonal or charge conjugated simple $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ minimal models. We apply the free-field resolution conjecture to Ishibashi states
$|L_{(\Lambda^{(+)},\Lambda^{(-)})}\rangle\rangle_{\Omega}=\sum_{\widehat{w}\in\widehat{W}}\theta_{\widehat{w}}|F_{(\widehat{w}\ast\widehat{\Lambda}^{(+)},\widehat{\Lambda}^{(-)})}\rangle\rangle_{\Omega^{\prime}},$ (29)
where the undetermined phase is there for the purpose of reproducing the correct chiral characters from the overlaps of free-field expressions. The form of $|F_{(\widehat{w}\ast\widehat{\Lambda}^{(+)},\widehat{\Lambda}^{(-)})}\rangle\rangle_{\Omega^{\prime}}$ is
$|F_{(\widehat{w}\ast\widehat{\Lambda}^{(+)},\widehat{\Lambda}^{(-)})}\rangle\rangle_{\Omega^{\prime}}=\exp{\Big{(}\frac{-1}{n}a_{-n}^{i}\cdot V_{\Omega^{\prime}}\cdot\widebar{a}_{-n}^{i}\Big{)}}\Big{|}p\otimes\widebar{p}\Big{\rangle}.$ (30)
We obtain the form of the gluing automorphism $V_{\Omega^{\prime}}$ by imposing the gluing conditions on the free-field Ishibashi states. Recall the form of free-field $\mathcal{W}\big{[}\widehat{g}(k)\big{]}$ energy-stress tensor (18), imposing the Virasoro gluing conditions leads to conditions similar to those of Kawai:2002vd ; Kawai:2002pz
$V_{\Omega^{\prime}}^{2}-I=0,\quad(\alpha_{+}\rho+\alpha_{-}\rho^{\vee})\cdot(V_{\Omega^{\prime}}-I)=0,$
$(p+\alpha_{+}\rho+\alpha_{-}\rho^{\vee})\cdot V_{\Omega^{\prime}}+\widebar{p}+(\alpha_{+}\rho+\alpha_{-}\rho^{\vee})=0.$ (31)
Next, we consider imposing the $\mathcal{W}$ gluing conditions. The $W^{i}(z)$ chiral generators other than the energy-stress tensor $T(z)$ are Virasoro primary operators. Hence, the following commutation relations hold
$L_{m},W^{i}_{n}]=\big{[}(h^{i}-1)m-n\big{]}W_{m+n}^{i}.$ (32)
Following the same calculation for the $\mathcal{W}_{3}$ case in Caldeira:2003zz , we write the $\mathcal{W}$ gluing conditions as
$\frac{1}{(h^{i}-1)n}(L_{n}-\widebar{L}_{-n})\big{[}W_{0}^{i}-(-1)^{h_{i}}\Omega\widebar{W}^{i}_{0}\big{]}|L_{(\Lambda^{(+)},\Lambda^{(-)})}\rangle\rangle_{\Omega}=0.$ (33)
(33) means that all the $W^{i}$ gluing conditions are satisfied if the Virasoro gluing condition, and the anti-holomorphic momentum satisfy the required conditions.
Next, we need to determine how the $W^{i}(z)$ eigenvalues $w^{i}$ change under the relation
$\widebar{p}=-p-2\alpha_{+}\rho-2\alpha_{-}\rho^{\vee}.$ (34)
We conjecture that $w^{i}$ eigenvalues are the standard symmetric polynomial of order $i$ of the variables $\theta_{j}=(\Lambda+\alpha_{+}\rho+\alpha_{-}\rho^{\vee},\epsilon_{j})$, where $\{\epsilon_{j}\}$ is the set of weight of vector representation of $g$. This conjecture is a direct generalization of the $w^{3}$ eigenvalue shown in Bouwknegt:1992wg . The $w^{i}$ eigenvalues changed under the action $\Lambda\to-\Lambda-\alpha_{+}\rho-\alpha_{-}\rho^{\vee}$ as $w^{i}\to(-1)^{h^{i}}w^{i}$. We show that this is how $w^{i}$ eigenvalues change under charge conjugation. The action of charge conjugation $U$ on the KM $\widehat{g}(k)$ currents as the conjugation automorphism $H^{i}(z)\to-H^{i}(z)$, $E^{\pm\alpha}(z)\to E^{\mp\alpha}(z)$, combined with the fact that $W^{i}(z)$ generators are obtained from the Hamiltonian reductions of the KM currents with $h^{i}$ order. Hence, the $w^{i}$ eigenvalues tramsform under charge conjugation as
$U:\quad w^{i}\to(-1)^{h^{i}}w^{i}.$ (35)
Hence, we can understand that the shift $\Lambda\to-\Lambda-\alpha_{+}\rho-\alpha_{-}\rho^{\vee}$ is a form of charge conjugation in the free-field realization.
From the above discussion, we conclude that the trivial $W^{i}_{0}$ gluing conditions
$(W^{i}_{0}-(-1)^{h^{i}}\widebar{W}^{i}_{0})|L_{(\Lambda^{(+)},\Lambda^{(-)})}\rangle\rangle_{I}=0,$ (36)
are satisfied by the free-field Ishibashi states with $V_{\Omega^{\prime}}=I$, and $\widebar{p}=-p-2\alpha_{+}\rho-2\alpha_{-}\rho^{\vee}$, for all $W^{i}(z)$ generators. The charge conjugated $W^{i}_{0}$ gluing conditions
$(W^{i}_{0}-\widebar{W}^{i}_{0})|L_{(\Lambda^{(+)},\Lambda^{(-)})}\rangle\rangle_{U}=0,$ (37)
are satisfied by the relation $V_{\Omega^{\prime}}=U$, and $\widebar{p}=-U\cdot(p+\alpha_{+}\rho+\alpha_{-}\rho^{\vee})-\alpha_{+}\rho-\alpha_{-}\rho^{\vee}=w_{l}\cdot(p+\alpha_{+}\rho+\alpha_{-}\rho^{\vee})-\alpha_{+}\rho-\alpha_{-}\rho^{\vee}$, because the $p-\widebar{p}$ relation is a Weyl symmetry identification (21), leaving the $w^{i}$ eigenvalues invariant.
Generalized Coulomb gas formalism.
The method we use to calculate the $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ correlation functions, is a direct generalization of the Coulomb gas formalism Dotsenko:1984nm ; Dotsenko:1984ad ; Dotsenko:1985hi , where the sphere correlation functions are calculated from free-field vertex operators with insertions of intertwiners to satisfy the neutrality condition. We prove the validity of this method by showing that the integral expressions obtained are the solutions to the Belavin-Polyakov-Zamolodchikov (BPZ) type differential equations when the neutrality conditions are satisfied Belavin:1984vu .
The BPZ-type differential equations are induced from inserting a singular vector into the chiral correlation function and applying the contour integral technique
$\langle\chi(z)\cdots\rangle=D\langle\mathcal{O}(z)\cdots\rangle=0.$ (38)
The chiral algebra intertwiners map $\mathcal{A}$ primary vectors to $\mathcal{A}$ primary vectors, hence, we could write any singular vector as $|\chi\rangle=Q^{\chi}|v\rangle$, where $Q^{\chi}$ is an intertwiner and $|v\rangle$ is a primary vector Thorn:1984sn ; Bouwknegt:1992wg . This allow us to replace the singular operator by a commutator $Q^{\chi},\mathcal{O}(z)]$, for some local operator $\mathcal{O}(z)$. Write the intertwiner as contour integral over a screening operator, we obtain
$\oint dw\>\langle s(w)\mathcal{O}(z)\cdots\rangle.$ (39)
When the neutrality condition is satisfied, the vertex operator correlation function transform covariantly under conformal transformations. We transform the screening operator to $w=\infty$, and the behavior of the whole vertex operator correlation function is $\sim w^{-2}$. Performing the contour integral at $w=\infty$, we arrive at the desired result 0.
With the validity of our method proven, we introduce the form of vertex operators. The free-field vertex operators that could recreate the desired fusion rules take the form of Bouwknegt:1990aa
$V_{\Lambda}=\>:e^{i(\alpha_{+}\Lambda^{(+)}+\alpha_{-}\Lambda^{(-)})\cdot\phi^{i}}:(z)\>P,$ (40)
where $P$ is some polynomials of the $\mathcal{W}$ intertwiners. In this work, we omit the $P$ part, and call the exponential part the free-field vertex operators $e^{i(\alpha_{+}\Lambda^{(+)}+\alpha_{-}\Lambda^{(-)})\cdot\phi^{i}}:(z)$. The dual vertex operators are also introduced, with the shift of the $g$ weight as $\Lambda\to-\Lambda-2\alpha_{+}\rho-2\alpha_{-}\rho^{\vee}$. Both the dual vertex operators and vertex operators with $U\Lambda$ will be used to represent charge conjugations in the free-field approach.
The application to disk correlation function calculations have been pioneered in Kawai:2002pz ; Caldeira:2003zz ; Hemming:2004dm . The free-field vertex operator realization of a $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ disk bulk $n$-point functions take the form of
$_{\Omega^{\prime}}\langle\langle L_{p}|\prod_{j=1}^{M}\oint s^{\pm}_{i}(z^{(j)})\prod_{\widebar{j}=1}^{\widebar{M}}\oint\widebar{s}^{\pm}_{i}(\widebar{j}^{(j)})\prod_{k=1}^{N}V_{\alpha_{k}}(z_{k})\prod_{\widebar{k}=1}^{\widebar{N}}\widebar{V}_{\widebar{\alpha}_{k}}(\widebar{z}_{k})|0\rangle$,
[…]
$_{\Omega^{\prime}}\langle\langle L_{p}|\prod_{i=1}^{N}V_{\alpha_{i}}(z_{i})\prod_{\widebar{i}=1}^{\widebar{N}}\widebar{V}_{\widebar{\alpha}_{i}}(\widebar{z}_{i})|0\rangle=\delta\widebar{\delta}\prod_{i<j}z_{ij}^{\alpha_{i}\cdot\alpha_{j}}\times\prod_{\widebar{i}<\widebar{j}}\widebar{z}_{\widebar{i}\widebar{j}}^{\widebar{\alpha}_{\widebar{i}}\cdot\widebar{\alpha}_{\widebar{j}}}\times\prod_{i,\widebar{j}}(1-z_{i}\widebar{z}_{\widebar{j}})^{\alpha_{i}\cdot V_{\Omega^{\prime}}\cdot\widebar{\alpha}_{\widebar{j}}}$. (41)
where the $\delta$ and $\widebar{\delta}$ represent the holomorphic and anti-holomorphic neutrality conditions, respectively.
Techniques in integrals.
We only consider calculations of disk two-point functions, and there are two crucial techniques in the integrations.
  • We send the position $z_{2}\to 0$, which simultaneously send $\widebar{z}_{2}\to\infty$ from the doubling trick on the disk. We perform an inversion coordinate transformation $\widebar{z}_{2}\to 0$ and absorb the Jacobian into the normalization.
  • Consider an integral with $n$ holomorphic screening operators insertions. We deformed the contours into the following integral along segments (with the phase factors absorbed into the normalization)
    $\oint dz^{(1)}\oint dz^{(2)}\cdots\oint dz^{(n)}\longrightarrow\int_{z_{2}}^{z_{1}}dz^{(1)}\int_{z_{2}}^{z^{(1)}}dz^{(2)}\cdots\int_{z_{2}}^{z^{(n-1)}}dz^{(n)}.$ (42)
With the deformation (42), we can apply the Euler-type integral expression and the Taylor expansions of the Lauricella’s hypergeometric function $F_{D}^{(n)}\Big{[}a,b_{1},\cdots,b_{n};c;z_{1},\cdots,z_{n}\Big{]}$, $z_{i}\in\mathbb{C}$ repeatedly to obtain the final result. The Taylor expansion of the $F_{D}^{(n)}$ functions at the origin of $\mathbb{C}^{n}$ is given by
$F_{D}^{(n)}\Big{[}a,b_{1},\cdots,b_{n};c;z_{1},\cdots,z_{n}\Big{]}=\sum_{k_{i}\in\mathbb{N}}\frac{(a)_{\sum_{i=1}^{n}k_{i}}\prod_{i=1}^{n}(b)_{k_{i}}}{(c)_{\sum_{i=1}^{n}k_{i}}}\prod_{i=1}^{n}\frac{z_{i}}{k_{i}!},$ (43)
which is convergent when $\max{|z_{i}|}<1$. $(a)_{k}=\Gamma(a+k)/\Gamma(a)$ is the Pochhammer symbol. The Euler-type of integral expression of the $F_{D}^{(n)}$ functions are given by
$F_{D}^{(n)}(a,b_{1},\cdots,b_{n};c;z_{1},\cdots,z_{n})=$
$\frac{\Gamma(c)}{\Gamma(a)\Gamma(c-a)}\int_{0}^{1}dt\>t^{a-1}(1-t)^{c-a-1}\prod_{i=1}^{n}(1-z_{i}t)^{-b_{i}},$ (44)
where $\text{Re}(c-a)>0$ and $\text{Re}(a)>0$. The integral is divergent if the previous two conditions are not satisfied.
Disk one-point functions.
Calculations of disk one-point function don’t require any insertions of $\mathcal{W}$ intertwiners. Consider a minimal model with diagonal and charge conjugation modular invariants. Disk one-point functions with Cardy boundary conditions are
$_{U\Omega}\langle b_{(\nu^{(+)},\nu^{(-)}})|\Phi_{(\Lambda^{(+)},\Lambda^{(-)})}^{\Omega}(z_{1},\widebar{z}_{1})|0\rangle=$
$\frac{S^{-1}_{(\nu^{(+)},\nu^{(-)})(\Lambda^{(+)},\Lambda^{(-)})}}{\sqrt{S_{0(\Lambda^{(+)},\Lambda^{(-)})}}}\>_{U\Omega}\langle\langle L_{(\Lambda^{(+)},\Lambda^{(-)})}|\Phi_{(\Lambda^{(+)},\Lambda^{(-)})}^{\Omega}(z_{1},\widebar{z}_{1})|0\rangle.$ (45)
The free-field realizations of the disk one-point functions in diagonal models are
$_{U}\langle\langle L_{(\Lambda^{(+)},\Lambda^{(-)})}|V_{(\Lambda^{(+)},\Lambda^{(-)})}(z_{1})\widebar{V}_{(w_{l}\ast\Lambda^{(+)},w_{l}\ast^{\vee}\Lambda^{(-)})}(\widebar{z}_{1})|0\rangle$
$(1-|z_{1}|^{2})^{-\Lambda\cdot w_{l}\cdot\big{[}w_{l}\cdot(\Lambda+\alpha_{+}\rho+\alpha_{-}\rho^{\vee})-(\alpha_{+}\rho+\alpha_{-}\rho^{\vee})\big{]}}.$ (46)
The free-field realizations of that in charge-conjugate models are
$_{I}\langle\langle L_{(\Lambda^{(+)},\Lambda^{(-)})}|V_{(\Lambda^{(+)},\Lambda^{(-)})}(z_{1})\widebar{V}_{(\Lambda^{(+)},\Lambda^{(-)})}^{{\dagger}}(\widebar{z}_{1})|0\rangle$
$(1-|z_{1}|^{2})^{-\Lambda\cdot(\Lambda+2\alpha_{+}\rho+2\alpha_{-}\rho^{\vee})}=(1-|z_{1}|^{2})^{-2h_{(\Lambda^{(+)},\Lambda^{(-)})}}.$ (47)
2.3 Disk two-point function calculations in some $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ minimal models
In this subsection, we show detailed examples of calculating the modular $S$ matrices, fusion rules, and the disk two-point functions using with Cardy boundary states, using free-field approaches. We will use the simplest solution to the neutrality conditions for our calculation. We only show the integral contributions from a single Ishibashi state, without writing the the final summation of each contribution.
2.3.1 Virasoro minimal models
The Virasoro algebra $\mathfrak{Vir}(p,p^{\prime})$ is the principal $\mathcal{W}$ algebra of $\widehat{A}_{1}(k)$, with $k+2=p/p^{\prime}$. The Virasoro minimal Kac-table is
$\big{\{}L_{(r,s)}\>|\>1\leq r\leq p-1,\quad 1\leq s\leq p^{\prime}-1\big{\}}.$ (48)
The labels are related to the Dynkin labels of $\widehat{A}_{1}$ highest-weight by $r=\Lambda^{(+)}_{1}+1$ and $s=\Lambda^{(-)}_{1}+1$. The Weyl symmetry on the Kac-table is $L_{(r,s)}\equiv L_{(p-r,p^{\prime}-s)}$. The $A_{1}$ Weyl group has only two elements $W=\{I,s_{1}=w_{l}\}$. Hence, Virasoro modular $S$ matrices have a simple expression
$S_{(r,s)(r^{\prime},s^{\prime})}=\sqrt{\frac{8}{pp^{\prime}}}(-1)^{(1+rs^{\prime}+sr^{\prime})}\sin{\Big{(}\pi\frac{p^{\prime}}{p}rr^{\prime}\Big{)}}\sin{\Big{(}\pi\frac{p}{p^{\prime}}ss^{\prime}\Big{)}}.$ (49)
The Virasoro minimal fusion rules are
$L_{(r_{1},s_{1})}\times L_{(r_{2},s_{2})}=\sum_{r_{3}=|r_{1}-r_{2}|+1}^{\min(r_{1}+r_{2}-1,2p-r_{1}-r_{2}-1)}\>\sum_{s_{3}=|s_{1}-s_{2}|+1}^{\min(s_{1}+s_{2}-1,2p^{\prime}-s_{1}-s_{2}-1)}L_{(r_{3},s_{3})},$ (50)
where the summation of $r_{3}$ and $s_{3}$ jumps by $2$ at each step. The complete Virasoro minimal modular invariants share the remarkable ADE structure as the simply-laced Lie algebras Cappelli:1986hf ; Cappelli:1987xt . We consider only diagonal Virasoro minimal models.
The Virasoro screening operators are
$s^{\pm}(z)=\>:e^{-i\alpha_{\pm}\alpha_{1}\cdot\phi}(z):\>=\>:e^{-i\sqrt{2}\alpha_{\pm}\phi}(z):\>.$ (51)
The free-field vertex operators are
$V_{(r,s)}(z)=\>:e^{i\sqrt{2}\alpha_{(r,s)}\phi}:(z),\quad\alpha_{(r,s)}=\frac{(r-1)\alpha_{+}+(s-1)\alpha_{-}}{2},$ (52)
and the dual vertex operators are
$V_{(r,s)}^{{\dagger}}=\>:e^{-i\sqrt{2}(\alpha_{0}+\alpha_{(r,s)})}:(z)\>.$ (53)
Disk correlation functions in the $\mathfrak{Vir}(5,2)$ minimal model.
The simplest non-trivial Virasoro minimal model is the non-unitary $\mathfrak{Vir}(5,2)$ model. The $\mathfrak{Vir}(5,2)$ Kac-table is
$\big{\{}L_{(1,1)}\equiv L_{(4,1)},\>L_{(2,1)}\equiv L_{(3,1)}\big{\}},$ (54)
whose primary conformal weights are $h_{(1,1)}=0$ and $h_{(2,1)}=-\frac{1}{5}$. The $\mathfrak{Vir}(5,2)$ modular $S$ matrix is
$S=\begin{pmatrix}-\sqrt{\frac{5+\sqrt{5}}{10}}&\sqrt{\frac{5-\sqrt{5}}{10}}\\ \sqrt{\frac{5-\sqrt{5}}{10}}&\sqrt{\frac{5+\sqrt{5}}{10}}\end{pmatrix}$. (55)
It is easy to verify that $S^{2}=C=I$. The non-obvious non-zero fusion rules are
$L_{(2,1)}\times L_{(2,1)}=L_{(1,1)}+L_{(2,1)}.$ (56)
Consider primary disk two-point functions $\langle b_{(r,s)}|\Phi_{(2,1)}(z_{1},\widebar{z}_{1})\Phi_{(2,1)}(z_{2},\widebar{z}_{2})|0\rangle$. It consists of two types contributions: $I_{(2,1)(2,1)}^{(1,1)}:=\langle\langle L_{(1,1)}|\Phi_{(2,1)}(z_{1},\widebar{z}_{1})\Phi_{(2,1)}(z_{2},\widebar{z}_{2})|0\rangle$ and $I_{(2,1)(2,1)}^{(2,1)}:=\langle\langle L_{(2,1)}|\Phi_{(2,1)}(z_{1},\widebar{z}_{1})\Phi_{(2,1)}(z_{2},\widebar{z}_{2})|0\rangle$. A free-field vertex operator realization of the contribution $I_{(2,1)(2,1)}^{(1,1)}$ is
$\oint dz^{(1)}\widebar{V}_{(2,1)}(\widebar{z}_{2})\widebar{V}_{(2,1)}^{{\dagger}}(\widebar{z}_{1})s^{+}(z^{(1)})V_{(2,1)}(z_{1})V_{(2,1)}(z_{2}),$ (57)
whose result is
$|z_{1}|^{\frac{4}{5}}(1-|z_{1}|^{2})^{\frac{2}{5}}\frac{\Gamma^{2}(\frac{3}{5})}{\Gamma(\frac{6}{5})}\>_{1}F_{2}\Big{[}\frac{3}{5},\frac{4}{5};\frac{6}{5};|z_{1}|^{2}\Big{]}.$
A free-field vertex operator realization of the contribution $I_{(2,1)(2,1)}^{(2,1)}$ is
$\oint d\widebar{z}^{(1)}\widebar{V}_{(2,1)}(\widebar{z}_{2})\widebar{V}_{(2,1)}^{{\dagger}}(\widebar{z}_{1})\widebar{s}^{+}(\widebar{z}^{(1)})V_{(2,1)}(z_{1})V_{(2,1)}(z_{2}),$ (58)
whose result is
$|z_{1}|^{\frac{2}{5}}(1-|z_{1}|^{2})^{\frac{2}{5}}\frac{\Gamma(\frac{3}{5})\Gamma(\frac{1}{5})}{\Gamma(\frac{4}{5})}\>_{1}F_{2}\Big{[}\frac{3}{5},\frac{2}{5};\frac{4}{5};|z_{1}|^{2}\Big{]}.$ (59)
Disk two-point functions in the Ising model $\mathfrak{Vir}(4,3)$.
The simplest unitary model is the Ising model $\mathfrak{Vir}(4,3)$, whose results are given in Kawai:2002vd ; Kawai:2002pz . We do the calculations independently. The Ising Kac-table is
$\big{\{}L_{(1,1)}\cong L_{(3,2)},\>L_{(1,2)}\cong L_{(3,1)},\>L_{(2,2)}\cong L_{(2,1)}\big{\}},$ (60)
whose primary conformal weights are $h_{(1,1)}=0$, $h_{(1,2)}=\frac{1}{2}$, and $h_{(2,2)}=\frac{1}{16}$. The Ising modular $S$ matrix is
$S=\frac{1}{2}\begin{pmatrix}1&1&\sqrt{2}\\ 1&1&-\sqrt{2}\\ \sqrt{2}&-\sqrt{2}&0\end{pmatrix}$. (61)
It can be easily verified that $S^{2}=C=I$. The non-obvious non-zero fusion rules are
$L_{(2,1)}\times L_{(2,1)}=L_{(1,1)}+L_{(3,1)},\quad L_{(2,1)}\times L_{(3,1)}=L_{(2,1)},\quad L_{(1,2)}\times L_{(1,2)}=L_{(1,1)}.$ (62)
We show the calculations of disk two-point functions $\langle b_{(r,s)}|\Phi_{(2,1)}(z_{1},\widebar{z}_{1})\Phi_{(2,1)}(z_{2},\widebar{z}_{2})|0\rangle$, which consist of two contributions $I_{(2,1)(2,1)}^{(1,1)}$ and $I_{(2,1)(2,1)}^{(3,1)}$. A free-field vertex operator realization of $I_{(2,1)(2,1)}^{(1,1)}$ is
$\oint dz^{(1)}\widebar{V}_{(2,1)}(\widebar{z}_{2})\widebar{V}_{(2,1)}^{{\dagger}}(\widebar{z}_{1})s^{+}(z^{(1)})V_{(2,1)}(z_{1})V_{(2,1)}(z_{2}),$ (63)
whose result is
$|z_{1}|^{-\frac{1}{4}}(1-|z_{1}|^{2})^{-\frac{1}{8}}\frac{\Gamma^{2}(\frac{1}{4})}{\Gamma(\frac{1}{2})}\>_{1}F_{2}\Big{[}\frac{1}{4},-\frac{1}{4};\frac{1}{2};|z_{1}|^{2}\Big{]}.$ (64)
A free-field vertex operator realization of the contribution $I_{(2,1)(2,1)}^{(3,1)}$
$\oint d\widebar{z}^{(1)}\>\widebar{V}_{(2,1)}(\widebar{z}_{2})\widebar{V}^{{\dagger}}_{(2,1)}(\widebar{z}_{1})\widebar{s}^{+}(\widebar{z}^{(1)})V_{(2,1)}(z_{1})V_{(2,1)}(z_{2}),$ (65)
whose result is
$|z_{1}|^{\frac{3}{4}}(1-|z_{1}|^{2})^{-\frac{1}{8}}\frac{\Gamma(\frac{1}{4})\Gamma(\frac{5}{4})}{\Gamma(\frac{3}{2})}\>_{1}F_{2}\Big{[}\frac{1}{4},\frac{3}{4};\frac{3}{2};|z_{1}|^{2}\Big{]}.$ (66)
The Virasoro minimal model $\mathfrak{Vir}(5,3)$.
The $\mathfrak{Vir}(5,3)$ Kac-table is
$\big{\{}L_{(1,1)}\cong L_{(4,2)},\>L_{(2,1)}\cong L_{(3,2)},\>L_{(4,1)}\cong L_{(1,2)},\>L_{(2,2)}\cong L_{(3,1)}\big{\}},$ (67)
whose primary conformal weights are $h_{(1,1)}=1$, $h_{(2,1)}=-1/20$, $h_{(4,1)}=3/4$, and $h_{(2,2)}=1/5$. The $\mathfrak{Vir}(5,3)$ modular $S$ matrix is
$S=\frac{1}{2}\begin{pmatrix}\sqrt{(1+\frac{1}{\sqrt{5}})}&\sqrt{(1-\frac{1}{\sqrt{5}})}&-\sqrt{(1+\frac{1}{\sqrt{5}})}&-\sqrt{(1-\frac{1}{\sqrt{5}})}\\ \sqrt{(1-\frac{1}{\sqrt{5}})}&\sqrt{(1+\frac{1}{\sqrt{5}})}&\sqrt{(1-\frac{1}{\sqrt{5}})}&\sqrt{(1+\frac{1}{\sqrt{5}})}\\ -\sqrt{(1+\frac{1}{\sqrt{5}})}&\sqrt{(1-\frac{1}{\sqrt{5}})}&-\sqrt{(1+\frac{1}{\sqrt{5}})}&\sqrt{(1-\frac{1}{\sqrt{5}})}\\ -\sqrt{(1-\frac{1}{\sqrt{5}})}&\sqrt{(1+\frac{1}{\sqrt{5}})}&\sqrt{(1-\frac{1}{\sqrt{5}})}&-\sqrt{(1+\frac{1}{\sqrt{5}})}\end{pmatrix}$. (68)
It is easy to verify that $S^{2}=C=I$. The $\mathfrak{Vir}(5,3)$ fusion rules are
$L_{(2,1)}\times L_{(2,1)}=L_{(1,1)}+L_{(3,1)},\quad L_{(2,1)}\times L_{(2,2)}=L_{(1,2)}+L_{(2,1)},$
$L_{(1,2)}\times L_{(1,2)}=L_{(1,1)},\quad L_{(2,2)}\times L_{(2,2)}=L_{(1,1)}+L_{(3,1)}.$ (69)
Consider disk correlation functions with Cardy boundary conditions $\langle b_{(r,s)}|\Phi_{(2,1)}(z_{1},\widebar{z}_{1})\Phi_{(2,1)}(z_{2},\widebar{z}_{2})|0\rangle$, which consists of two contributions $I_{(2,1)(2,1)}^{(1,1)}$ and $I_{(2,1)(2,1)}^{(3,1)}$. A free-field vertex operator realization of contribution $I_{(2,1)(2,1)}^{(1,1)}$ is
$\oint dz^{(1)}\widebar{V}_{(2,1)}(\widebar{z}_{2})\widebar{V}_{(2,1)}^{{\dagger}}(\widebar{z}_{1})s^{+}(z^{(1)})V_{(2,1)}(z_{1})V_{(2,1)}(z_{2}),$ (70)
whose result is
$|z_{1}|^{\frac{1}{5}}(1-|z_{1}|^{2})^{\frac{1}{10}}\frac{\Gamma^{2}(\frac{2}{5})}{\Gamma(\frac{4}{5})}\>_{1}F_{2}\Big{[}\frac{2}{5},\frac{1}{5};\frac{4}{5};|z_{1}|^{2}\Big{]}.$ (71)
A free-field vertex operator realization of contribution $I_{(2,1)(2,1)}^{(3,1)}$ is
$\oint d\widebar{z}^{(1)}\widebar{V}_{(2,1)}(\widebar{z}_{2})\widebar{V}_{(2,1)}^{{\dagger}}(\widebar{z}_{1})\widebar{s}^{+}(\widebar{z}^{(1)})V_{(2,1)}(z_{1})V_{(2,1)}(z_{2}),$ (72)
whose result is
$|z_{1}|^{\frac{3}{5}}(1-|z_{1}|^{2})^{\frac{1}{10}}\frac{\Gamma(\frac{2}{5})\Gamma(\frac{4}{5})}{\Gamma(\frac{6}{5})}\>_{1}F_{2}\Big{[}\frac{2}{5},\frac{3}{5};\frac{6}{5};|z_{1}|^{2}\Big{]}.$ (73)
2.3.2 $\mathcal{W}_{3}(p,p^{\prime})$ minimal models
The $\mathcal{W}_{3}(p,p^{\prime})$ algebra is the principal $\mathcal{W}\big{[}\widehat{A}_{2}(k)\big{]}$ algebra, with $k=p/p^{\prime}-3$. The Cartan matrix of $A_{2}$ is
$A=\begin{pmatrix}2&-1\\ -1&2\end{pmatrix}$. (74)
The $A_{2}$ simple roots are $\alpha_{1}=[2,-1]$ and $\alpha_{2}=[-1,2]$. The Weyl group $W$ has $6$ elements, generated by the simple Weyl reflections $w_{1}=\begin{pmatrix}-1&0\\ 1&1\end{pmatrix}$ and $w_{2}=\begin{pmatrix}1&1\\ 0&-1\end{pmatrix}$. The longest element has length $l(w_{l})=3$. The $A_{2}$ highest root is $\theta=\alpha_{1}+\alpha_{2}=\alpha_{1}^{\vee}+\alpha_{2}^{\vee}$. The $\widehat{A}_{2}$ Coxeter and the dual Coxeter numbers are $h=h^{\vee}=3$. The $\mathcal{W}_{3}(p,p^{\prime})$ minimal Kac-table is
$\widehat{\Lambda}^{(+)}\in\widehat{P}_{+}^{p-3},\quad\widehat{\Lambda}^{(-)}\in\widehat{P}_{+}^{p^{\prime}-3}.$ (75)
The $\widehat{A}_{2}$ Weyl symmetries permute the Dynkin labels of $\widehat{\Lambda}^{(\pm)}$.
Disk two-point functions in $\mathcal{W}_{3}(5,3)$ model.
The simplest non-unitary $\mathcal{W}_{3}$ minimal model is the $\mathcal{W}_{3}(5,3)$ model. The only allowed $\Lambda^{(-)}$ is $[0,0]$, and there are a total of six choice of $\Lambda^{(+)}$. We label the $\mathcal{W}_{3}(5,3)$ irreducible modules by $\Lambda^{(+)}$ only, denoted by $L_{\Lambda^{(+)}}$. The $\mathcal{W}_{3}(5,3)$ minimal Kac-table is
$\Big{\{}L_{[0,0]}\equiv L_{[2,0]}\equiv L_{[0,2]},\>L_{[1,1]}\equiv L_{[1,0]}\equiv L_{[0,1]}\Big{\}}.$ (76)
Both the irreducible modules are self charge-conjugated. The primary conformal weights of the two modules are
$h_{[0,0]}=0,\quad h_{[1,1]}=-\frac{1}{5}.$ (77)
The $\mathcal{W}_{3}(5,3)$ modular $S$ matrix is
$S=\begin{pmatrix}-\sqrt{\frac{1}{2}(1+\frac{1}{\sqrt{5}})}&\sqrt{\frac{1}{2}(1-\frac{1}{\sqrt{5}})}\\ \sqrt{\frac{1}{2}(1-\frac{1}{\sqrt{5}})}&\sqrt{\frac{1}{2}(1+\frac{1}{\sqrt{5}})}\end{pmatrix}$. (78)
It is easy to verify that $S^{2}=C=I$. The non-obvious non-zero fusion rules are given by
$L_{[1,1]}\times L_{[1,1]}=L_{[0,0]}+L_{[1,1]}.$ (79)
Consider the disk two-point functions with Cardy boundary conditions $\langle b_{\Lambda^{(+)}}|\Phi_{[1,1]}(z_{1},\widebar{z}_{1})\Phi_{[1,1]}(z_{2},\widebar{z}_{2})|0\rangle$ in the diagonal $\mathcal{W}_{3}(5,3)$ model, which consists of two contributions $I_{[1,1][1,1]}^{[0,0]}$ and $I_{[1,1][1,1]}^{[1,1]}$. A free-field vertex operators realization of $I_{[1,1][1,1]}^{[0,0]}$ is
$\oint dz^{(1)}\oint dz^{(2)}\widebar{V}_{[1,1]}(\widebar{z}_{2})\widebar{V}^{{\dagger}}_{[1,1]}(\widebar{z}_{1})s_{1}^{+}(z^{(1)})s_{2}^{+}(z^{(2)})V_{[1,0]}(z_{1})V_{[0,1]}(z_{2}),$ (80)
whose result is
$|z_{1}|^{\frac{4}{5}}(1-|z_{1}|^{2})^{\frac{1}{5}}\frac{\Gamma^{2}(\frac{2}{5})}{\Gamma(\frac{4}{5})}\sum_{k\in\mathbb{N}}\frac{(\frac{1}{5})_{k}(\frac{2}{5})_{k}}{k!(\frac{4}{5})_{k}}\frac{\Gamma(\frac{2}{5})\Gamma(\frac{4}{5}+k)}{\Gamma(\frac{6}{5}+k)}\>_{1}F_{2}\Big{[}k+\frac{4}{5},\frac{1}{5};k+\frac{6}{5};|z_{1}|^{2}\Big{]}.$ (81)
A free-field vertex operator realization of the contribution is $I_{[1,1][1,1]}^{[1,1]}$
$\oint d\widebar{z}^{(1)}\oint d\widebar{z}^{(2)}\widebar{V}_{[1,1]}(\widebar{z}_{2})\widebar{V}_{[1,1]}^{{\dagger}}(\widebar{z}_{1})\widebar{s}^{+}_{1}(\widebar{z}^{(1)})\widebar{s}^{+}_{2}(\widebar{z}^{(2)})V_{[1,0]}(z_{1})V_{[0,1]}(z_{2}),$ (82)
whose results is
$|z_{1}|^{\frac{2}{5}}(1-|z_{1}|^{2})^{\frac{1}{5}}\frac{\Gamma^{2}(\frac{2}{5})}{\Gamma(\frac{4}{5})}\sum_{k\in\mathbb{N}}\frac{(\frac{1}{5})_{k}(\frac{2}{5})_{k}}{k!(\frac{4}{5})_{k}}\frac{\Gamma(\frac{1}{5}+k)\Gamma(\frac{4}{5})}{\Gamma(1+k)}\>_{1}F_{2}\Big{[}\frac{1}{5}+k,\frac{1}{5};1+k;|z_{1}|^{2}\Big{]}.$ (83)
The $\mathcal{W}_{3}(5,4)$ minimal models.
The simplest unitary $\mathcal{W}_{3}$ minimal model is the $\mathcal{W}_{3}(5,4)$ minimal model. The diagonal $\mathcal{W}_{3}(5,4)$ minimal model contains a non-diagonal Virasoro sub-theory, which is the three-state Potts model. Studying the three-state Potts model using free-field approaches to $\mathcal{W}_{3}$ minimal models has been given in Caldeira:2003zz . We consider the free-field approach to disk two-point functions of the diagonal and the charge-conjugated $\mathcal{W}_{3}(5,4)$ minimal models. The $\mathcal{W}_{3}(5,4)$ Kac-table is
$\Lambda^{(+)}\in\big{\{}[0,0],[2,0],[0,2],[1,0],[1,1],[0,1]\big{\}},\quad\Lambda^{(-)}\in\big{\{}[0,0],[1,0],[0,1]\big{\}}.$ (84)
The $\mathbb{Z}_{3}$ Weyl symmetry allows us to fix the choice of $\Lambda^{(-)}=[0,0]$, and label all irreducible modules by the six $\Lambda^{(+)}$. The modules $L_{[0,0]}$ and $L_{[1,1]}$ are self-conjugated, while $UL_{[2,0]}=L_{[0,2]}$ and $UL_{[1,0]}=L_{[0,1]}$. The primary conformal weights of them are
$h_{[0,0]}=0,\>h_{[1,1]}=\frac{2}{5},\>h_{[2,0]}=h_{[0,2]}=\frac{2}{3},\>h_{[1,0]}=h_{[0,1]}=\frac{1}{15},$ (85)
The modular $S$ matrix of this model is given by
$S=\frac{1}{10\sqrt{3}}\begin{pmatrix}c&d&c&c&d&d\\ d&-c&d&d&-c&-c\\ c&d&a&a^{\ast}&b&b^{\ast}\\ c&d&a^{\ast}&a&b^{\ast}&b\\ d&-c&b&b^{\ast}&-a&-a^{\ast}\\ d&-c&b^{\ast}&b&-a^{\ast}&-a\end{pmatrix}$, (86)
where $a=2i(e^{\frac{2\pi i}{15}}-2e^{-\frac{4\pi i}{15}}-e^{\frac{8\pi i}{15}}+2e^{\frac{14\pi i}{15}})$, $b=2i(2e^{-\frac{2\pi i}{15}}+e^{\frac{4\pi i}{15}}-2e^{-\frac{8\pi i}{15}}-e^{-\frac{14\pi i}{15}})$, $c=\sqrt{50-10\sqrt{5}}$, and $d=\sqrt{50+10\sqrt{5}}$. It is easy to verify that $S^{2}=C$. From which we can obtain six Cardy boundary states for both diagonal and charge conjugate models. Applying the Verlinde formula, we obtain the non-obvious non-zero fusion rules
$L_{[1,1]}\times L_{[1,1]}=L_{[0,0]}+L_{[1,1]},\quad L_{[0,1]}\times L_{[0,1]}=L_{[1,0]},\quad L_{[1,0]}\times L_{[1,0]}=L_{[0,1]},$
$L_{[2,0]}\times L_{[0,2]}=L_{[0,0]}+L_{[1,1]},\quad L_{[0,2]}\times L_{[0,2]}=L_{[2,0]},\quad L_{[2,0]}\times L_{[2,0]}=L_{[0,2]}.$ (87)
We begin with the example $\langle b_{\Lambda^{(+)}}|\Phi_{[1,0]}(z_{1},\widebar{z}_{1})\Phi_{[1,0]}(z_{2},\widebar{z}_{2})|0\rangle$ which is different in diagonal and charge conjugated models. In the charge conjugated model, a free-field realization of the contribution $I_{[1,0][1,0]}^{[0,1]}$ is
$\oint d\widebar{z}^{(1)}\oint dz^{(1)}\widebar{V}_{[1,0]}^{{\dagger}}(\widebar{z}_{2})\widebar{V}_{[0,1]}(\widebar{z}_{1})\widebar{s}_{2}^{+}(\widebar{z}^{(1)})s_{1}^{+}(z^{(1)})V_{[1,0]}(z_{1})V_{[1,0]}(z_{2}),$ (88)
whose result is
$\sum_{k\in\mathbb{N}}|z_{1}|^{-\frac{2}{15}+2k}(1-|z_{1}|^{2})^{\frac{4}{15}}\frac{(\frac{1}{5})_{k}(\frac{4}{5})_{k}}{k!(\frac{2}{5})_{k}}\frac{\Gamma^{2}(\frac{1}{5})}{\Gamma(\frac{2}{5})}\frac{\Gamma(\frac{1}{5})\Gamma(\frac{3}{5}+k)}{\Gamma(\frac{4}{5}+k)}.$ (89)
In the diagonal model, a free-field realization of the contribution $I_{[1,0][1,0]}^{[0,1]}$ is
$\oint d\widebar{z}^{(1)}\oint dz^{(1)}\widebar{V}_{[0,1]}^{{\dagger}}(\widebar{z}_{2})\widebar{V}_{[1,0]}(\widebar{z}_{1})\widebar{s}_{1}^{+}(\widebar{z}^{(1)})s_{1}^{+}(z^{(1)})V_{[1,0]}(z_{1})V_{[1,0]}(z_{2}),$ (90)
whose result is
$\sum_{k}|z_{1}|^{-\frac{2}{15}+2(k_{1}+k_{2})}(1-|z_{1}|^{2})^{\frac{4}{15}}\frac{\Gamma^{2}(\frac{1}{5})}{\Gamma(\frac{2}{5})}\frac{(\frac{1}{5})_{\sum k_{i}}(-\frac{8}{5})_{k_{1}}(\frac{4}{5})_{k_{2}}}{k_{1}!k_{2}!(\frac{2}{5})_{\sum k_{i}}}\frac{\Gamma(\frac{3}{5})\Gamma(\frac{1}{5})}{\Gamma(\frac{4}{5})}\>_{1}F_{2}\Big{[}\frac{1}{5},\frac{4}{5};\frac{3}{5};|z_{1}|^{2}\Big{]}.$ (91)
The two different results reflects the difference between charge conjugated and diagonal models.
Next, we consider the disk correlation function $\langle b_{\Lambda^{(+)}}|\Phi_{[1,1]}(z_{1},\widebar{z}_{1})\Phi_{[1,1]}(z_{2},\widebar{z}_{2})|0\rangle$, whose results in the diagonal and charge conjugate models are equal. The contribution $I_{[1,1][1,1]}^{[0,0]}$ has a free-field vertex operator realization
$\oint dz^{(1)}\oint dz^{(2)}\oint dz^{(3)}\oint dz^{(4)}\widebar{V}_{[1,1]}(\widebar{z}_{2})\widebar{V}_{[1,1]}^{{\dagger}}(\widebar{z}_{1})$
$s^{+}_{2}(z^{(4)})s^{+}_{2}(z^{(3)})s^{+}_{1}(z^{(2)})s^{+}_{1}(z^{(1)})V_{[1,1]}(z_{1})V_{[1,1]}(z_{2}),$ (92)
whose result is
$\sum_{k,l,m}|z_{1}|^{-\frac{8}{5}+2k_{4}+2l_{3}+2m_{2}}(1-|z_{1}|^{2})^{\frac{8}{5}}\frac{(\frac{1}{5})_{\sum k_{i}}(\frac{4}{5})_{k_{1}}(\frac{4}{5})_{k_{2}}(\frac{4}{5})_{k_{3}}(-\frac{2}{5})_{k_{4}}}{(\frac{14}{5})_{\sum k_{i}}k_{1}!\cdots k_{4}!}$
$\frac{(\sum k_{i}+2)_{\sum_{l_{i}}}(\frac{4}{5})_{l_{1}}(\frac{4}{5})_{l_{2}}(-\frac{2}{5})_{l_{3}}}{(\sum k_{i}+\frac{11}{5})_{\sum_{l_{i}}}l_{1}!l_{2}!l_{3}!}\frac{(a^{\prime})_{\sum m_{i}}(\frac{4}{5})_{m_{1}}(-\frac{2}{5})_{m_{1}}}{(c^{\prime})_{\sum m_{i}}m_{1}!m_{2}!}$
$\frac{\Gamma(\frac{1}{5})\Gamma(\sum_{i}k_{i}+2)\Gamma(a^{\prime})\Gamma(\frac{13}{5})\Gamma(a)\Gamma(\frac{4}{5})}{\Gamma(\sum_{i}k_{i}+\frac{11}{5})\Gamma(c^{\prime})\Gamma(c)}\>_{1}F_{2}\Big{[}a,-\frac{2}{5};c;|z_{1}|^{2}\Big{]},$ (93)
where $a^{\prime}=\frac{3}{5}+k_{1}+k_{3}+k_{4}+\sum l_{i}$, $c^{\prime}=\frac{16}{5}+k_{1}+k_{3}+k_{4}+\sum l_{i}$, $a=\frac{4}{5}+k_{3}+k_{4}+l_{2}+l_{3}+m_{2}$, and $c=\frac{8}{5}+k_{3}+k_{4}+l_{2}+l_{3}+m_{2}$. A free-field realization of the contribution $I_{[1,1][1,1]}^{[1,1]}$ is
$\oint dz^{(1)}\oint dz^{(2)}\oint d\widebar{z}^{(1)}\oint d\widebar{z}^{(2)}\widebar{V}^{{\dagger}}_{[1,1]}(\widebar{z}_{2})\widebar{V}_{[1,1]}(\widebar{z}_{1})$
$\widebar{s}^{+}_{1}(\widebar{z}^{(1)})\widebar{s}^{+}_{2}(\widebar{z}^{(2)})s^{+}_{1}(z^{(1)})s^{+}_{2}(z^{(2)})V_{[1,1]}(z_{1})V_{[1,1]}(z_{2}).$ (94)
The finite part of this integral is
$\sum_{k,m,l}|z_{1}|^{-\frac{8}{5}+2k_{4}+2m_{3}+2l_{2}}(1-|z_{1}|^{2})^{\frac{4}{5}}\frac{(\frac{1}{5})_{\sum k_{i}}(\frac{4}{5})_{k_{1}}(\frac{4}{5})_{k_{2}}(\frac{4}{5})_{k_{3}}(\frac{4}{5})_{k_{4}}}{k_{1}!k_{2}!k_{3}!k_{4}!(\frac{2}{5})_{\sum k_{i}}}$
$\frac{\Gamma^{2}(\frac{1}{5})}{\Gamma(\frac{2}{5})}\frac{\Gamma(\frac{1}{5})\Gamma(-\frac{2}{5}+\sum k_{i})}{\Gamma(-\frac{1}{5}+\sum k_{i})}\frac{(-\frac{2}{5}+\sum k_{i})_{\sum l_{i}}(-\frac{8}{5})_{l_{1}}(\frac{4}{5})_{l_{2}}(\frac{4}{5})_{l_{3}}}{l_{1}!l_{2}!l_{3}!(-\frac{1}{5}+\sum k_{i})_{\sum l_{i}}}$
$\frac{\Gamma(\frac{1}{5})\Gamma(a^{\prime})}{\Gamma(c^{\prime})}\frac{(a^{\prime})_{\sum m_{i}}(\frac{4}{5})_{m_{1}}(\frac{4}{5})_{m_{2}}}{(c^{\prime})_{\sum m_{i}}}\frac{\Gamma(\frac{1}{5})\Gamma(a)}{\Gamma(c)}\>_{1}F_{2}\Big{[}a,\frac{4}{5};c;|z_{1}|^{2}\Big{]}.$ (95)
The parameters are $a^{\prime}=\frac{7}{5}+l_{2}+k_{3}$, $c^{\prime}=\frac{8}{5}+l_{2}+k_{3}$, $a=2+k_{2}+k_{3}+l_{1}+l_{2}+m_{1}+m_{2}$, and $c=\frac{11}{5}+k_{2}+k_{3}+l_{1}+l_{2}+m_{1}+m_{2}$. The divergent part contains the integral
$\int dt\>t^{-\frac{7}{5}}(1-t)^{-\frac{4}{5}}(1-tz_{1}\widebar{z}^{(1)})^{\frac{8}{5}}(1-tz_{1}\widebar{z}^{(2)})^{\frac{8}{5}}(1-t|z_{1}|^{2})^{-\frac{4}{5}}.$ (96)
2.3.3 The $\mathcal{W}_{4}(7,4)$ minimal model
The $\mathcal{W}_{4}(p,p^{\prime})$ algebra is the principal $\mathcal{W}\big{[}\widehat{A}_{3}(k)\big{]}$ algebra at level $k=\frac{p}{p^{\prime}}-4$. The $A_{3}$ Cartan matrix is
$A=\begin{pmatrix}2&-1&0\\ -1&2&-1\\ 0&-1&2\end{pmatrix}$, (97)
with the simple roots and coroots being $\alpha_{1}=\alpha_{1}^{\vee}=[2,-1,0]$, $\alpha_{2}=\alpha_{2}^{\vee}=[-1,2,-1]$, and $\alpha_{3}=\alpha_{3}^{\vee}=[0,-1,2]$. The $A_{3}$ Weyl group has a total of $24$ elements, generated by three simple Weyl reflections $w_{1}=\begin{pmatrix}-1&0&0\\ 1&1&0\\ 0&0&1\end{pmatrix}$, $w_{2}=\begin{pmatrix}1&1&0\\ 0&-1&0\\ 0&1&1\end{pmatrix}$, and $w_{3}=\begin{pmatrix}1&0&0\\ 0&1&1\\ 0&0&-1\end{pmatrix}$. The longest element has length $l(w_{l})=6$. The highest root of $A_{3}$ is $\theta=\alpha_{1}+\alpha_{2}+\alpha_{3}=\alpha_{1}^{\vee}+\alpha_{2}^{\vee}+\alpha_{3}^{\vee}$, hence the Coxeter and the dual Coxeter numbers of $\widehat{A}_{3}$ are $h=h^{\vee}=4$. $|P/Q^{\vee}|=\det A_{ij}=4$. The Weyl symmetry acting on the Kac-table as a $\mathbb{Z}_{4}$ permutation group on $\widehat{\Lambda}^{(\pm)}$. The non-degenerate $\mathcal{W}_{4}(7,4)$ minimal Kac-table is
$\Big{\{}L_{[0,0,0]},\quad L_{[0,1,0]},\quad L_{[1,1,1]},\quad L_{[1,0,0]},\quad L_{[0,0,1]}\Big{\}},$ (98)
with $L_{[0,0,0]}$, $L_{[0,1,0]}$, and $L_{[1,1,1]}$ being self charge conjugated modules, and $UL_{[1,0,0]}=L_{[0,0,1]}$. The primary conformal weights of them are
$h_{[0,0,0]}=0,\quad h_{[0,1,0]}=-\frac{4}{7},\quad h_{[1,1,1]}=-\frac{5}{7},\quad h_{[1,0,0]}=h_{[0,0,1]}=-\frac{3}{7}.$ (99)
The $\mathcal{W}_{4}(7,4)$ minimal modular $S$ matrix is
$S=\frac{1}{\sqrt{7}}\begin{pmatrix}-\frac{2}{7}a&-\frac{2}{7}b&-\frac{2}{7}c&1&1\\ -\frac{2}{7}b&-\frac{2}{7}c&\frac{2}{7}a&-1&-1\\ -\frac{2}{7}c&\frac{2}{7}a&\frac{2}{7}b&1&1\\ 1&-1&1&d&d^{\ast}\\ 1&-1&1&d^{\ast}&d\end{pmatrix}$, (100)
where $a=(-1+5\cos{\frac{2\pi}{7}}-2\cos{\frac{4\pi}{7}}-2\cos{\frac{6\pi}{7}})$, $b=(1+2\cos{\frac{2\pi}{7}}+2\cos{\frac{4\pi}{7}}-5\cos{\frac{6\pi}{7}})$, , $c=(-1-2\cos{\frac{2\pi}{7}}+5\cos{\frac{4\pi}{7}}-2\cos{\frac{6\pi}{7}})$, and $d=\frac{1}{2}i(i+\sqrt{7})$. It is easy to verify that $S^{2}=C$. The Verlinde formula provides the following non-obvious fusion rules
$L_{[0,1,0]}\times L_{[0,1,0]}=L_{[0,0,0]}+L_{[0,1,0]}+L_{[1,1,1]},$
$L_{[0,1,0]}\times L_{[1,1,1]}=L_{[0,1,0]}+L_{[1,1,1]}+L_{[1,0,0]}+L_{[0,0,1]}$
$L_{[1,1,1]}\times L_{[1,1,1]}=L_{[0,0,0]}+L_{[0,1,0]}+2L_{[1,1,1]}+L_{[1,0,0]}+L_{[0,0,1]},$
$L_{[1,0,0]}\times L_{[1,0,0]}=L_{[0,0,1]},\quad L_{[0,0,1]}\times L_{[0,0,1]}=L_{[1,0,0]},$
$L_{[1,0,0]}\times L_{[0,0,1]}=L_{[0,0,0]}+L_{[1,1,1]}.$ (101)
Consider the disk two-point functions $\langle b_{\Lambda^{(+)}}|\Phi_{[0,1,0]}(z_{1},\widebar{z}_{1})\Phi_{[0,1,0]}(z_{2},\widebar{z}_{2})|0\rangle$, with a total of three contributions $I_{[0,1,0][0,1,0]}^{[0,0,0]}$, $I_{[0,1,0][0,1,0]}^{[0,1,0]}$, and $I_{[0,1,0][0,1,0]}^{[1,1,1]}$. A free-field vertex operator realization of $I_{[0,1,0][0,1,0]}^{[0,0,0]}$ is
$\oint dz^{(1)}\oint dz^{(2)}\oint dz^{(3)}\oint dz^{(4)}\widebar{V}^{{\dagger}}_{[0,1,0]}(\widebar{z}_{2})\widebar{V}_{[0,1,0]}(\widebar{z}_{1})$
$s_{3}^{+}(z^{(4)})s_{1}^{+}(z^{(3)})s_{2}^{+}(z^{(2)})s_{2}^{+}(z^{(1)})V_{[0,1,0]}(z_{1})V_{[0,1,0]}(z_{2}),$ (102)
whose the result is
$\sum_{k,l,m}|z_{1}|^{\frac{16}{7}+2m_{2}}(1-|z_{1}|^{2})^{\frac{4}{7}}\frac{(1)_{\sum k}(\frac{4}{7})_{k_{1}}(\frac{4}{7})_{k_{2}}}{k_{1}!k_{2}!(2)_{\sum k}}\frac{\Gamma(\frac{3}{7})}{\Gamma(\frac{17}{7})}\frac{(2)_{l}(\frac{4}{7})_{l}}{l!(\frac{17}{7})_{l}}$
$\frac{\Gamma(a^{\prime})\Gamma(\frac{15}{7})}{\Gamma(c^{\prime})}\frac{(a^{\prime})_{\sum m_{i}}(\frac{4}{7})_{m_{1}}(\frac{4}{7})_{m_{2}}}{m_{1}!m_{2}!(c^{\prime})_{\sum m_{i}}}\frac{\Gamma(a)\Gamma(\frac{3}{7})}{\Gamma(c)}\>_{1}F_{2}\Big{[}a,\frac{4}{7};c;|z_{1}|^{2}\Big{]},$
where $k_{i},l,m_{i}\in\mathbb{N}$, $a^{\prime}=\frac{9}{7}+k_{2}+l$, $c^{\prime}=\frac{24}{7}+k_{2}+l$, $a=\frac{12}{7}+\sum m_{i}$, and $c=\frac{15}{7}+\sum m_{i}$. A free-field vertex operator realization of $I_{[0,1,0][0,1,0]}^{[0,1,0]}$ is
$\oint d\widebar{z}^{(1)}\oint d\widebar{z}^{(2)}\oint d\widebar{z}^{(3)}\oint d\widebar{z}^{(4)}\widebar{V}_{[0,1,0]}^{{\dagger}}(\widebar{z}_{2})\widebar{V}_{[0,1,0]}(\widebar{z}_{1})$
$\widebar{s}_{3}^{+}(\widebar{z}^{(4)})\widebar{s}_{1}^{+}(\widebar{z}^{(3)})\widebar{s}_{2}^{+}(\widebar{z}^{(2)})\widebar{s}_{2}^{+}(\widebar{z}^{(1)})V_{[0,1,0]}(z_{1})V_{[0,1,0]}(z_{2}),$ (103)
whose result includes one divergent term, we write down the finite part
$\sum_{k,l,m}|z_{1}|^{\frac{8}{7}+2m_{2}}(1-|z_{1}|^{2})^{\frac{4}{7}}\frac{\Gamma(\frac{1}{7})}{\Gamma(\frac{8}{7})}\frac{(\frac{1}{7})_{\sum k_{i}}(\frac{4}{7})_{k_{1}}(\frac{4}{7})_{k_{2}}}{k_{1}!k_{2}!\frac{8}{7})_{\sum k_{i}}}\frac{\Gamma(a^{\prime})\Gamma(\frac{3}{7})}{\Gamma(c^{\prime})}\frac{(a^{\prime})_{l}(\frac{4}{7})_{l}}{l!(c^{\prime})_{l}}\frac{\Gamma(a)\Gamma(\frac{15}{7})}{\Gamma(c)}$
$\frac{(a)_{\sum m_{i}}(\frac{4}{7})_{m_{1}}(\frac{4}{7})_{m_{2}}}{m_{1}!m_{2}!(c)_{\sum m_{i}}}\frac{\Gamma(\frac{4}{7}+\sum m_{i})\Gamma(\frac{3}{7})}{\Gamma(1+\sum m_{i})}\>_{1}F_{2}\Big{[}\frac{4}{7}+\sum m_{i},\frac{4}{7};1+\sum m_{i};|z_{1}|^{2}\Big{]}.$ (104)
The parameters are $a^{\prime}=\frac{2}{7}+\sum k_{i}$ and $c^{\prime}=\frac{5}{7}+\sum k_{i}$, $a=-\frac{1}{7}+k_{2}+l$, $c=2+k_{2}+l$, $k_{2}+l\in\mathbb{Z}_{+}$ and $k_{2},l\in\mathbb{N}$. The divergent term contains the integral
$\int_{0}^{1}dt\>t^{-\frac{8}{7}}(1-t)^{\frac{8}{7}}\Big{[}1-t\frac{\widebar{z}^{(1)}}{\widebar{z}_{1}}\Big{]}^{-\frac{4}{7}}(1-t\widebar{z}^{(1)}z_{1})^{-\frac{4}{7}}.$ (105)
The final contribution $I_{[0,1,0][0,1,0]}^{[1,1,1]}$ has a free-field realization as
$\oint d\widebar{z}^{(3)}\oint d\widebar{z}^{(2)}\oint d\widebar{z}^{(1)}\oint dz^{(1)}\widebar{V}^{{\dagger}}_{[0,1,0]}(\widebar{z}_{2})\widebar{V}_{[0,1,0]}(\widebar{z}_{1})$
$\widebar{s}_{3}^{+}(\widebar{z}^{(1)})\widebar{s}_{1}^{+}(\widebar{z}^{(2)})\widebar{s}_{2}^{+}(\widebar{z}^{(3)})s_{2}^{+}(z^{(1)})V_{[0,1,0]}(z_{1})V_{[0,1,0]}(z_{2})$ (106)
The finite part is
$\sum_{k,l,m}|z_{1}|^{\frac{6}{7}+2k_{3}+2k_{4}+2l+2m}(1-|z_{1}|^{2})^{\frac{4}{7}}\frac{\Gamma(\frac{5}{7})\Gamma(\frac{3}{7})}{\Gamma(\frac{8}{7})}\frac{(\frac{5}{7})_{\sum k_{i}}(\frac{4}{7})_{k_{1}}(\frac{4}{7})_{k_{2}}(-\frac{8}{7})_{k_{3}}(\frac{4}{7})_{k_{4}}}{k_{1}!k_{2}!k_{3}!k_{4}!(\frac{8}{7})_{\sum k_{i}}}$
$\frac{\Gamma(\frac{2}{7}+\sum k_{i})}{\Gamma(\frac{9}{7}+\sum k_{i})}\frac{(\frac{2}{7}+\sum k_{i})_{l}(\frac{4}{7})_{l}}{l!(\frac{9}{7}+\sum k_{i})_{l}}\frac{\Gamma(a^{\prime})}{\Gamma(c^{\prime})}\frac{(a^{\prime})_{m}(\frac{4}{7})_{m}}{m!(c^{\prime})_{m}}\frac{\Gamma(\frac{3}{7})\Gamma(a)}{\Gamma(c)}\>_{1}F_{2}\Big{[}a,\frac{4}{7},c;|z_{1}|^{2}\Big{]}.$ (107)
The parameters $a^{\prime}=-\frac{1}{7}+k_{2}+k_{3}+k_{4}+l$ and $c^{\prime}=\frac{6}{7}+k_{2}+k_{3}+k_{4}+l$, $a^{\prime}>0$. The divergent term contains the integral
$\int_{0}^{1}dt\>t^{-\frac{8}{7}}(1-t)^{0}(1-tz^{(1)}\widebar{z}_{1})^{-\frac{4}{7}}.$ (108)
2.3.4 The $\mathcal{W}\big{[}\widehat{G}_{2}\big{]}(6,7)$ minimal model
We consider one principal QDS $\mathcal{W}$ minimal model related to the non-simply-laced Lie algebra, which is the $\mathcal{W}\big{[}\widehat{G}_{2}\big{]}(6,7)$ model. The $G_{2}$ Cartan matrix is
$A=\begin{pmatrix}2&-3\\ -1&2\end{pmatrix}$. (109)
The simple roots and coroots are $\alpha_{1}=\alpha_{1}^{\vee}=[2,-3]$ and $\alpha_{2}=\frac{1}{3}\alpha_{2}^{\vee}=[-1,2]$. Hence, $r^{\vee}=3$, and the $G_{2}$ Weyl vector and dual Weyl vector are $\rho=[1,1]$ and $\rho^{\vee}=[1,3]$. The highest root $\theta=2\alpha_{1}+3\alpha_{2}=2\alpha_{1}+\alpha_{2}^{\vee}$. The Coxeter and the dual of Coxeter number of $\widehat{G}_{2}$ are $h=6$ and $h^{\vee}=4$, respectively. $|P/Q^{\vee}|=\det(\alpha_{i}^{\vee},\alpha_{j}^{\vee})=3$. The $G_{2}$ Weyl group has a total of 12 elements, generated by the two simple Weyl reflections $w_{1}=\begin{pmatrix}-1&0\\ 3&1\end{pmatrix}$ and $w_{2}=\begin{pmatrix}1&1\\ 0&-1\end{pmatrix}$. The longest Weyl element has length $l(w_{l})=6$. The zeroth Dynkin label of a $\widehat{G}_{2}$ weight vector $\widehat{\lambda}$ is given by
$\lambda_{0}=k-2\lambda_{1}-\lambda_{2}.$ (110)
Hence, the Kac table of the $\mathcal{W}\big{[}\widehat{G}_{2}\big{]}(6,7)$ minimal model is
$\Lambda^{(+)}\in\big{\{}[0,0],[1,0],[0,1],[0,2]\big{\}},\quad\Lambda^{(-)}=[0,0].$ (111)
We label them using $\Lambda^{(+)}$, and their primary conformal weights are given by
$h_{[0,0]}=0,\quad h_{[1,0]}=-\frac{1}{3},\quad h_{[0,1]}=-\frac{2}{3},\quad h_{[0,2]}=-\frac{5}{9}.$ (112)
The $\mathcal{W}\big{[}\widehat{G}_{2}\big{]}(6,7)$ modular $S$ matrix is
$S=\frac{1}{3\sqrt{3}}\begin{pmatrix}a&b&c&-3\\ b&-c&-a&-3\\ c&-a&b&3\\ -3&-3&3&0\end{pmatrix}$, (113)
$a=(\cos{\frac{8\pi}{9}}+\cos{\frac{5\pi}{9}}-\cos{\frac{4\pi}{9}}-\cos{\frac{\pi}{9}})$, $b=(\cos{\frac{\pi}{9}}+\cos{\frac{2\pi}{9}}-\cos{\frac{7\pi}{9}}-\cos{\frac{8\pi}{9}})$, $c=(\cos{\frac{2\pi}{9}}-\cos{\frac{4\pi}{9}}+\cos{\frac{5\pi}{9}}-\cos{\frac{7\pi}{9}})$. It is easy to verify that $S^{2}=C=I$. Applying the Verlinde formula to this model, we obtain the non-obvious non-zero fusion rules
$L_{[1,0]}\times L_{[1,0]}=L_{[0,0]}+L_{[0,2]},\quad L_{[1,0]}\times L_{[0,1]}=L_{[0,1]}+L_{[0,2]},$
$L_{[0,1]}\times L_{[0,1]}=L_{[0,0]}+L_{[1,0]}+L_{[0,1]}+L_{[0,2]},\quad L_{[0,2]}\times L_{[0,2]}=L_{[0,0]}+L_{[0,1]}.$ (114)
Consider the diak two-point functions $\langle b_{\Lambda^{(+)}}|\Phi_{[1,0]}(z_{1},\widebar{z}_{1})\Phi_{[0,1]}(z_{2},\widebar{z}_{2})|0\rangle$. The contribution $I_{[1,0][0,1]}^{[0,1]}$ is realized by
$\oint d\widebar{z}^{(1)}\oint dz^{(5)}\oint dz^{(4)}\oint dz^{(3)}\oint dz^{(2)}\oint dz^{(1)}\widebar{V}_{[0,1]}(\widebar{z}_{2})\widebar{V}_{[1,0]}^{{\dagger}}(\widebar{z}_{1})\widebar{s}_{2}^{+}(\widebar{z}^{(1)})$
$s^{+}_{1}(z^{(5)})s^{+}_{1}(z^{(2)})s^{+}_{2}(z^{(3)})s^{+}_{2}(z^{(4)})s^{+}_{2}(z^{(5)})V_{[1,0]}(z_{1})V_{[0,1]}(z_{2}).$
Unlike the previous examples, where we can separate some of the divergent terms and obtain a finite part. The $\widebar{z}^{(1)}$ integral is divergent, indicate all terms in this expression has a divergent overall multiplier. The divergent problem might be solved by choosing a different solution to the neutrality conditions. However, other solutions to the neutrality conditions are significantly more complicated. Hence, we write down only the integral expression, without performing the explicit integral. A free-field realization of $I_{[1,0][0,1]}^{[0,2]}$ is
$\oint d\widebar{z}^{(4)}\oint d\widebar{z}^{(3)}\oint d\widebar{z}^{(2)}\oint d\widebar{z}^{(1)}\oint dz^{(2)}\oint dz^{(1)}\widebar{V}_{[0,1]}(\widebar{z}_{2})\widebar{V}^{{\dagger}}_{[1,0]}(\widebar{z}_{1})$
$\widebar{s}_{1}^{+}(\widebar{z}^{(4)})\widebar{s}_{2}^{+}(\widebar{z}^{(3)})\widebar{s}_{2}^{+}(\widebar{z}^{(2)})\widebar{s}_{2}^{+}(\widebar{z}^{(1)})s^{+}_{2}(z^{(2)})s^{+}_{1}(z^{(1)})V_{[1,0]}(z_{1})V_{[0,1]}(z_{2}),$ (115)
which has a similar overall divergent factor like the contribution $I_{[1,0][0,1]}^{[0,1]}$.
Appendix A The background charged bosonic free-field approach to semi-simple bosonic QDS $\mathcal{W}$ minimal models
We discuss the rational principle QDS $\mathcal{W}$ minimal models related to bosonic semi-simple KM algebras Kac:1984mq ; Kac:1990gs . Our discussion focuses on the free-field approach to them, that is, constructing the $\mathcal{W}$ algebra as the centralizer of the screening operators.
Consider a KM algebra $\widehat{g}$ related to a semi-simple Lie algebra $g=g_{1}\oplus g_{2}\oplus\cdots\oplus g_{n}$. Each simple part $\widehat{g}_{i}$ has a central extension, denoted by $k^{(i)}$, and its Coxeter number and dual Coxeter number, denoted as $h^{(i)}$ and $h^{\vee(i)}$ respectively. Analogous to the simple cases, we define parameters $(\alpha_{+}^{(i)})^{-2}=k^{(i)}+h^{\vee(i)}$ and $\alpha_{+}^{(i)}\alpha_{-}^{(i)}=-1$. We denote the simple roots as
$\Big{\{}\{\alpha_{i}^{(1)}\},\{\alpha_{i}^{(2)}\},\cdots,\{\alpha_{i}^{(n)}\}\Big{\}},$ (116)
where $i=1,2,\cdots,r_{1}+r_{2}+\cdots+r_{n}$, and the superscript labels the corresponding simple parts. The simple roots from different simple parts are in different dimensions in the weight space. The simple coroots and fundamental weights are defined in the same manner as in the simple cases. A semi-simple weight vector $\Lambda$ is denoted by
$\Lambda=\big{[}\Lambda_{1},\Lambda_{2},\cdots,\Lambda_{r_{1}+\cdots+r_{n}}\big{]}=\big{(}\Lambda^{(1)},\Lambda^{(2)},\cdots,\Lambda^{(n)}\big{)}.$ (117)
The Wakimoto-type of free-field realization generalize to semi-simple cases straight-forwardly. Hence, the $\mathcal{W}$ screening operators induced from the QDS procedure take the same form as the simple cases $e^{-i\alpha_{\pm}\alpha^{(j)}_{i}\cdot\phi}:(z)$. The chiral generators of the semi-simple QDS $\mathcal{W}$ algebra are given by
$\Big{\{}\big{[}T^{(1)}(z),\mathcal{W}^{i(1)},\cdots\big{]},\cdots,\big{[}T^{(n)}(z),\mathcal{W}^{i(n)},\cdots\big{]}\Big{\}},$ (118)
where each block $\big{[}T^{(a)}(z),\mathcal{W}^{i(a)},\cdots\big{]}$ consist of the free-field simple QDS $\mathcal{W}\big{[}\widehat{g}_{a}(k^{(a)})\big{]}$ chiral generators. The commutation relations satisfy
$\mathcal{W}^{i(a)}_{m},\mathcal{W}^{j(b)}_{n}]\quad\propto\quad\delta^{ab}.$ (119)
A semi-simple $\mathcal{W}\big{[}\widehat{g}(k^{(1)},\cdots,k^{(n)})\big{]}$ Verma module $V_{\Lambda}$ has a primary vector $|v_{\Lambda}\rangle$ satisfying
$W^{i(a)}_{n>0}|v_{\Lambda}\rangle=0,\quad W^{i(a)}_{0}|v_{\Lambda}\rangle=w^{i(a)}|v_{\Lambda}\rangle,\quad\forall i,a\>.$ (120)
The semi-simple $\mathcal{W}$ Verma modules are tensor products of simple $\mathcal{W}$ Verma modules
$V_{\Lambda}=V_{\Lambda^{(1)}}^{\widehat{g}_{1}}\otimes\cdots\otimes V_{\Lambda^{(n)}}^{\widehat{g}_{n}}.$ (121)
The singular vector structure of $V_{\Lambda}$ is the combination of all the simple $V_{\Lambda^{(i)}}^{\widehat{g}_{i}}$ singular vector structures. The chiral characters of the semi-simple $\mathcal{W}$ modules are the products of simple $\mathcal{W}$ chiral characters
$\text{ch}_{\Lambda}=\text{ch}_{\Lambda^{(1)}}^{\mathcal{W}[\widehat{g}_{1}]}\times\text{ch}_{\Lambda^{(2)}}^{\mathcal{W}[\widehat{g}_{2}]}\times\cdots\times\text{ch}_{\Lambda^{(n)}}^{\mathcal{W}[\widehat{g}_{n}]}.$ (122)
(122) indicates that the semi-simple $\mathcal{W}$ modular matrices are products of the simple $\mathcal{W}$ modular matrices
$T_{\Lambda\mu}=\prod_{i=1}^{n}T_{\Lambda^{(i)}\mu^{(i)}},\quad S_{\Lambda\mu}=\prod_{i=1}^{n}S_{\Lambda^{(i)}\mu^{(i)}}.$ (123)
A semi-simple $\mathcal{W}\big{[}\widehat{g}(k^{(1)},k^{(2)},\cdots,k^{(n)})\big{]}$ algebra is rational when all simple $\mathcal{W}\big{[}\widehat{g}_{j}(k^{(j)})\big{]}$ are rational. Hence, the semi-simple minimal Kac-table is the combination of all simple Kac-tables
$\widehat{\Lambda}^{(+i)}\in\widehat{P}_{+}^{p^{i}-h^{\vee i}},\quad\widehat{\Lambda}^{(-i)}\in\widehat{P}_{+}^{\vee\>(p^{\prime})^{i}-h^{i}}.$ (124)
The Weyl group of a semi-simple Lie algebra is the direct product of the simple Weyl groups
$W(g)\cong W(g_{1})\times W(g_{2})\times\cdots\times W(g_{n}),$ (125)
hence the expression for the simple $\mathcal{W}$ minimal modular matrix (22) is valid for the semi-simple cases. We apply the Verlinde formula to the rational semi-simple $\mathcal{W}$ minimal models. The relation (123) reveals that the fusion rules of the semi-simple model are the product of the simple $\mathcal{W}$ minimal fusion rules
$N_{\Lambda\mu}^{\nu}=\prod_{i=1}^{n}N_{\Lambda^{(i)}\mu^{(i)}}^{\nu^{(i)}}.$ (126)
The Fock space vertex operators in the semi-simple $\mathcal{W}$ algebras are
$e^{i\Lambda\cdot\phi}:(z)=\>:e^{i\Lambda^{(1)}\cdot\phi^{(1)}}:(z)\times:e^{i\Lambda^{(2)}\cdot\phi^{(2)}}:(z)\times\cdots\times:e^{i\Lambda^{(n)}\cdot\phi^{(n)}}:(z).$ (127)
The Fock space primary is obtained by
$|v_{\Lambda}\rangle=:e^{i\Lambda\cdot\phi}:(z)|0\rangle=|v_{\Lambda^{(1)}}\rangle\otimes\cdots\otimes|v_{\Lambda^{(n)}}\rangle.$ (128)
The free-field resolution of an irreducible minimal module is an $n$-dimensional complex $(F_{\widehat{\Lambda}},\>d^{(1)},\cdots,d^{(n)})$, where $d^{(a)}$ are simple $\mathcal{W}\big{[}\widehat{g}_{a}(k^{(a)})\big{]}$ intertwiners. The free-field vertex operators approach to correlation function calculations is analogous to the Coulomb gas formalism to the simple cases. The correlation functions of the semi-simple $\mathcal{W}$ minimal models are again the multiplication of the simple $\mathcal{W}$ minimal parts. This can be shown directly from the integral expression, since the integrals factorize multiplications of independent integrals over simple $\mathcal{W}$ minimal models.
Appendix B A Free-field approach to coset $\mathcal{W}$ minimal Ishibashi states
We show a free-field approach to principal ADE $\mathcal{W}$ coset algebra minimal Ishibashi states. Consider a coset CFT $\widehat{g}(k)/\widehat{h}(jk)$, $h\subset g$. Its energy-stress tensor is given by
$T(z)=T^{\widehat{g}(k)}(z)-T^{\widehat{h}(jk)}(z).$ (129)
All the coset operators are singular operators of $\widehat{h}(jk)$. A $\widehat{g}(k)$ module $L_{\widehat{\Lambda}}^{\widehat{g}}$ decomposes as the finite direct sum
$L_{\widehat{\Lambda}}^{\widehat{g}}=\bigoplus_{\widehat{\lambda}}\big{(}L_{(\widehat{\Lambda},\widehat{\lambda})}^{\widehat{g}/\widehat{h}}\otimes L_{\widehat{\lambda}}^{\widehat{h}}\big{)}$, (130)
where $L_{(\widehat{\Lambda},\widehat{\lambda})}^{\widehat{g}/\widehat{h}}$ are coset modules. Taking the trace over the (130), we obtain
$\text{ch}_{\widehat{\Lambda}}=\sum_{\widehat{\lambda}}b_{(\widehat{\Lambda},\widehat{\lambda})}\times\text{ch}_{\widehat{\lambda}}$, (131)
where the coset chiral characters are defined to be the branching functions $b_{(\widehat{\Lambda},\widehat{\lambda})}$.
We state the logic of obtaining the coset resolution complex. Assume that a free-field resolution complex of the $\widehat{g}$ module $L_{\widehat{\Lambda}}$ is $(F_{\widehat{\Lambda}},d)$. The vectors in the coset modules $L_{(\widehat{\Lambda},\widehat{\lambda})}^{\widehat{g}/\widehat{h}}$ are $\widehat{h}$ singular vectors with fixed $\widehat{h}$ weight $\widehat{\lambda}$. Hence, the subcomplex $(S_{(\widehat{\Lambda},\widehat{\lambda})},d)$ of $(F_{\widehat{\Lambda}},d)$ is the resolution complex of coset module $L^{\widehat{g}/\widehat{h}}_{(\widehat{\Lambda},\widehat{\lambda})}$, where $S_{(\widehat{\Lambda},\widehat{\lambda})}\subset F_{\widehat{\Lambda}}$ consists of $\widehat{h}$ singular vectors with fixed $\widehat{\lambda}$ Bouwknegt:1990fb ; Bouwknegt:1990wa . The universal method of this projection is the BRST procedure, where we project to desired $\widehat{h}-$weight using a $\text{rank}(h)$-dimensional constraint condition. Hence, $\text{rank}(h)$ bcbc ghost systems are introduced.
The coset ADE principal $\mathcal{W}\big{[}\widehat{g}\big{]}$ theories are realized by the diagonal cosets Goddard:1984vk ; Goddard:1986ee ; Mathieu:1990dy ; Bouwknegt:1992wg
$\widehat{g}^{(1)}(k)\oplus\widehat{g}^{(2)}(1)/\widehat{g}^{(3)}(k+1),$ (132)
where diagonal means that the KM currents in the three parts satisfy the diagonal embedding $J^{a(3)}(z)=J^{a(1)}(z)+J^{a(2)}(z)$. We denote the diagonal cosets by $\big{[}\widehat{g}\oplus\widehat{g}/\widehat{g},(k,1)\big{]}$. The coset chiral characters of the ADE diagonal cosets $\big{[}\widehat{g}\oplus\widehat{g}/\widehat{g},(k,1)\big{]}$ are identical to that of QDS $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ algebras, at $k+h^{\vee}=p/(p^{\prime}-p):=p/u$ Bouwknegt:1992wg . A rigorous proof of the equivalence between simply-laced principal QDS $\mathcal{W}$ algebras and the diagonal coset $\mathcal{W}$ algebra was given in Arakawa:2018iyk . The ADE $\big{[}\widehat{g}\oplus\widehat{g}/\widehat{g},(k,1)\big{]}$ branching rules are Bouwknegt:1992wg
$L^{(1)}_{\widehat{\Lambda}^{(1)}}\otimes L^{(2)}_{\widehat{\Lambda}^{(2)}}=\bigoplus\>\big{(}L^{\mathcal{W}}_{(\Lambda^{(+)},\Lambda^{(-)})}\otimes L^{(3)}_{\widehat{\Lambda}^{(3)}}\big{)},$ (133)
where $\Lambda^{(1)}+\Lambda^{(2)}-\Lambda^{(3)}\in Q$, $\widehat{\Lambda}^{(1)}=\widehat{\Lambda}^{(+)}-(u-1)(k^{\prime}+h^{\vee})\omega_{0}$ $\widehat{\Lambda}^{(2)}\in\widehat{P}_{+}^{1}$, and $\widehat{\Lambda}^{(3)}=\widehat{\Lambda}^{(-)}-(u-1)(k^{\prime}+h^{\vee}+1)\omega_{0}$ ($\omega_{0}$ being the zeroth fundamental weight).
The coset resolution has been applied to diagonal ADE coset modules in Bouwknegt:1990fb . The projection is realized by the BRST approach, where we introduce $r=\text{rank}(g)$ bcbc ghost systems. The BRST invariant projection operator is
$\widehat{P}_{\Lambda^{(3)}}=\int_{0}^{1}d\theta\>e^{2\pi i\theta\cdot(\widehat{h}_{0}-{\Lambda}^{(3)})}$ (134)
where $\widehat{h}_{0}$ is the BRST invariant ($r$-dimensional) Cartan current
$\widehat{h}_{0}=h_{0}^{(1)}+h_{0}^{(2)}+\text{ghost terms}.$ (135)
The resolution complex of the tensor products of the two dominant modules $L_{\widehat{\Lambda}^{(1)}}\otimes L_{\widehat{\Lambda}^{(2)}}$ is the double complex $(\widetilde{F}_{\widehat{\Lambda}^{(1)}}\otimes\widetilde{F}_{\widehat{\Lambda}^{(2)}},\>\widetilde{d}^{(1)},\>\widetilde{d}^{(2)})$, where $\widetilde{d}^{(1)},\widetilde{d}^{(2)}]=0$. Applying the $H_{Q}\big{(}F^{bc}\otimes F^{\beta\gamma}\big{)}\cong\mathbb{C}$ lemma to the $\widetilde{F}_{\widehat{\Lambda}^{(1)}}\otimes\widetilde{F}_{\widehat{\Lambda}^{(2)}}\otimes F^{bc}$ Fock space modules, we obtain that the Fock space modules in the projected complex are
$F_{\widehat{w}^{(1)}\ast\widehat{\Lambda}^{(1)}}^{\phi^{i}}\otimes F_{\widehat{w}^{(2)}\ast\widehat{\Lambda}^{(2)}}^{\phi^{i}}\otimes F^{\beta^{\alpha}\gamma^{\alpha}}.$ (136)
The form of the resolution complex are then
$\Big{[}\widehat{P}_{w^{(1)}\ast\Lambda^{(1)}+w^{(2)}\ast\Lambda^{(2)}-\Lambda^{(3)}}(F_{\widehat{w}^{(1)}\ast\widehat{\Lambda}^{(1)}}^{\phi}\otimes F_{\widehat{w}^{(2)}\ast\widehat{\Lambda}^{(2)}}^{\phi}\otimes F^{\beta^{\alpha}\gamma^{\alpha}})\>,\widetilde{d}^{(1)},\>\widetilde{d}^{(2)}\Big{]},$ (137)
where the zeroth cohomology space of it is conjectured to be isomorphic to the $\mathcal{W}$ irreducible modules $L_{(\Lambda^{(+)},\Lambda^{(-)})}$.
The trivial and charge conjugated gluing conditions of three $\widehat{g}$ algebras transform into the that in the coset model. The trivial and charge conjugated $\widehat{g}$ gluing conditions also transform into that of the $\phi^{i}\beta^{\alpha}\gamma^{\alpha}$ systems. Hence, by applying the free-field resolution conjectures to the coset $\mathcal{W}$ Ishibashi states, we obtain
$|L_{(\Lambda^{(+)},\Lambda^{(-)})}\rangle\rangle_{\Omega}=\sum_{\widehat{w}^{(1,2)}}\widehat{P}_{w^{(1)}\ast\Lambda^{(1)}+w^{(2)}\ast\Lambda^{(2)}-\Lambda^{(3)}}\theta_{\widehat{w}_{1},\widehat{w}_{2}}$
$|F_{\widehat{w}^{(1)}\ast\widehat{\Lambda}^{(1)}}\rangle\rangle_{\Omega}\otimes|F_{\widehat{w}^{(2)}\ast\widehat{\Lambda}^{(2)}}\rangle\rangle_{\Omega}\otimes|F^{\beta^{\alpha}\gamma^{\alpha}}\rangle\rangle_{I},$ (138)
where $\Omega=I,U$, and the phases $\theta_{\widehat{w}_{1},\widehat{w}_{2}}$ are introduced to produce the correct phase in the overlap. The free-field Ishibashi states $|F_{\widehat{w}^{(i)}\ast\widehat{\Lambda}^{(i)}}\rangle\rangle_{\Omega}$, $i=1,2$, take the form of
$|F_{\widehat{w}^{(i)}\ast\widehat{\Lambda}^{(i)}}\rangle\rangle_{\Omega}=\exp{\Big{[}\sum_{n\in\mathbb{Z}_{+}}\frac{-1}{n}a_{-n}\cdot V_{\Omega}\cdot\widebar{a}_{-n}\Big{]}}|p\otimes\widebar{p}\rangle,\quad p=\alpha_{+}\widehat{w}^{(i)}\ast\widehat{\Lambda}^{(i)}.$ (139)
The analysis of the form of the $V_{\Omega}$ and the $p-\widebar{p}$ relation is exactly the same as the $\mathcal{W}\big{[}\widehat{g}(k)\big{]}$ cases. Hence for the diagonal and conjugated cases, $V_{\Omega}=I,U$, and $\widebar{p}=-V_{\Omega}\cdot(p+\alpha_{+}\rho)-\alpha_{+}\rho$. The $\beta^{\alpha}\gamma^{\alpha}$ Ishibashi states take the form of
$|F^{\beta^{\alpha}\gamma^{\alpha}}\rangle\rangle_{I}=\exp{\Big{[}-\sum_{\alpha\in\Delta_{+}}\sum_{n\in\mathbb{Z}_{+}}(\beta_{-n}\widebar{\gamma}_{-n}+\gamma_{-n}\widebar{\beta}_{-n})\Big{]}}|0\otimes\widebar{0}\rangle.$ (140)
Acknowledgements.
The author would like to thank H.Z.Liang for guidance and encouragement.